L^{A}TEX: Structured documents for TEX

unofficial L^{A}TEX reference manual version 2.0, June 2008 This document is version 2.0 (June 2008) of an unofficial reference manual for L^{A}TEX, a document preparation system. L^{A}TEX is a macro package for TEX. This was originally translated from ‘LATEX.HLP’ v1.0a in the VMS Help Library. The pretranslation version was written by George D. Greenwade of Sam Houston State University. The L^{A}TEX 2.09 version was written by Stephen Gilmore. The L^{A}TEX2e version was adapted from this by Torsten Martinsen. Karl Berry made further updates and additions, and gratefully acknowledges using Hypertext Help with L^{A}TEX, by Sheldon Green, and the L^{A}TEX Command Summary (for L^{A}TEX 2.09) by L. Botway and C. Biemesderfer (published by the TEX Users Group as TEXniques number 10), as reference material (text was not directly copied). Copyright \begin{matrix}c \\ ◯ \end{matrix} 2007, 2008 Karl Berry. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988, 1994, 2007 Stephen Gilmore. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1994, 1995, 1996 Torsten Martinsen. Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies. Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one. Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions.

Short Contents

LATEX2e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1 Overview of LATEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2 Starting & ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 3 Document classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4 Typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 5 Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 6 Sectioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 7 Cross references. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 8 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 9 Footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 10 Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 11 Counters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 12 Lengths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 13 Line breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 14 Page breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 15 Making paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 16 Math formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 17 Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 18 Page Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 19 Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 20 Boxes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 21 Special insertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 22 Splitting the input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 23 Front/back matter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 24 Letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 25 Terminal Input/Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 26 Command Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 A Document templates. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 Command Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

Table of Contents

L^{A}TEX2e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

1 Overview of L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2 Starting & ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

3 Document classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

3.1 Document class options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4 Typefaces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

4.1 Font styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 4.2 Font sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 4.3 Low-level font commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

5 Layout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

5.1 \onecolumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.2 \twocolumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 5.3 \flushbottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.4 \raggedbottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.5 Page layout parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11

6 Sectioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

7 Cross references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

7.1 \label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.2 \pageref{key} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 7.3 \ref{key}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13

8 Environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

8.1 abstract . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.2 array. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 8.3 center . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

8.3.1 \centering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8.4 description. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8.5 displaymath. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 8.6 document . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 8.7 enumerate . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 8.8 eqnarray . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8.9 equation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8.10 figure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 8.11 flushleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

8.11.1 \raggedright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 8.12 flushright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

8.12.1 \raggedleft. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.13 itemize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 8.14 letter environment: writing letters . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.15 list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.16 math. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.17 minipage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 8.18 picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

8.18.1 \circle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8.18.2 \makebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8.18.3 \framebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8.18.4 \dashbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8.18.5 \frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.6 \line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.7 \linethickness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.8 \thicklines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.9 \thinlines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.10 \multiput . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.11 \oval. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 8.18.12 \put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 8.18.13 \shortstack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 8.18.14 \vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

8.19 quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 8.20 quote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 8.21 tabbing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 8.22 table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 8.23 tabular . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

8.23.1 \multicolumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 8.23.2 \cline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.23.3 \hline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.23.4 \vline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

8.24 thebibliography. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.24.1 \bibitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 8.24.2 \cite. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.24.3 \nocite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.24.4 Using BibTEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

8.25 theorem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 8.26 titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 8.27 verbatim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

8.27.1 \verb. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 8.28 verse. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

9 Footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

9.1 \footnote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.2 \footnotemark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.3 \footnotetext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 9.4 Footnote parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

10 Definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35

10.1 \newcommand & \renewcommand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.2 \newcounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.3 \newlength. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 10.4 \newsavebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36

10.5 \newenvironment & \renewenvironment . . . . . . . . . . . . . . . . . . . . 36

10.6 \newtheorem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 10.7 \newfont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 10.8 \protect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

11 Counters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38

11.1 \alph \Alph \arabic \roman \Roman \fnsymbol: Printing

counters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 11.2 \usecounter{counter}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 11.3 \value{counter} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 11.4 \setcounter{counter}{value} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.5 \addtocounter{counter}{value}. . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.6 \refstepcounter{counter} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.7 \stepcounter{counter} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 11.8 \day \month \year: Predefined counters . . . . . . . . . . . . . . . . . . . . . 39

12 Lengths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

12.1 \setlength{\len}{value} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12.2 \addtolength{\len}{amount} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12.3 \settodepth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12.4 \settoheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12.5 \settowidth{\len}{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 12.6 Predefined lengths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40

13 Line breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41

13.1 \\[*][morespace] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.2 \obeycr & \restorecr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.3 \newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.4 \- (discretionary hyphen) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.5 \fussy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.6 \sloppy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 13.7 \hyphenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 13.8 \linebreak & \nolinebreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

14 Page breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

14.1 \cleardoublepage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 14.2 \clearpage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 14.3 \newpage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 14.4 \enlargethispage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 14.5 \pagebreak & \nopagebreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43

15 Making paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . 44

15.1 \indent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 15.2 \noindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 15.3 \parskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 15.4 Marginal notes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44

16 Math formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

16.1 Subscripts & Superscripts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 16.2 Math symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 16.3 Math functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 16.4 Math accents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 16.5 Spacing in math mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 16.6 Math Miscellany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

17 Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

18 Page Styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

18.1 \maketitle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 18.2 \pagenumbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 18.3 \pagestyle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 18.4 \thispagestyle{style} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60

19 Spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61

19.1 \hspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 19.2 \hfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 19.3 \SPACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 19.4 \@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 19.5 \thinspace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 19.6 \/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.7 \hrulefill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.8 \dotfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.9 \addvspace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.10 \bigskip \medskip \smallskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.11 \vfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 19.12 \vspace[*]{length} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63

20 Boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

20.1 \mbox{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 20.2 \fbox and \framebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 20.3 lrbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 20.4 \makebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 20.5 \parbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 20.6 \raisebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 20.7 \savebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 20.8 \sbox{\boxcmd}{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 20.9 \usebox{\boxcmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66

21 Special insertions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

21.1 Reserved characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 21.2 Text symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 21.3 Accents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 21.4 Non-English characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 21.5 \rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 21.6 \today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69

22 Splitting the input . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

22.1 \include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 22.2 \includeonly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 22.3 \input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70

23 Front/back matter . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

23.1 Tables of contents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 23.1.1 \addcontentsline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 23.1.2 \addtocontents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71

23.2 Glossaries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 23.3 Indexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

24 Letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

24.1 \address{return-address} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 24.2 \cc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 24.3 \closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 24.4 \encl. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.5 \location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.6 \makelabels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.7 \name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.8 \opening{text}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.9 \ps . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.10 \signature{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.11 \startbreaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.12 \stopbreaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 24.13 \telephone . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75

25 Terminal Input/Output. . . . . . . . . . . . . . . . . . . . . . 76

25.1 \typein[cmd]{msg} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 25.2 \typeout{msg} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76

26 Command Line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

Appendix A Document templates . . . . . . . . . . . . . 78

A.1 book template . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 A.2 beamer template. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 A.3 tugboat template . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79

Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81

Command Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

L^{A}TEX2e

This document is version 2.0 (June 2008) of an unofficial reference manual for L^{A}TEX, a document preparation system. It is intended to cover L^{A}TEX2e, which has been the standard version of L^{A}TEX for many years.

L^{A}TEX is implemented as a macro package for Donald E. Knuth’s TEX typesetting program. L^{A}TEX was originally created by Leslie Lamport; it is now maintained by a group of volunteers (http://latex-project.org). The official documentation written by the L^{A}TEX project is available there. Again, the present document is unofficial and has not been reviewed by the L^{A}TEX maintainers.

1 Overview of L^{A}TEX

The L^{A}TEX command typesets a file of text using the TEX program and the L^{A}TEX “macro package” for TEX. To be more specific, it processes an input file containing the text of a document with interspersed commands that describe how the text should be formatted. It produces at least three files as output:

1. A main output file, which is one of:

1. If invoked as latex, a “Device Independent” (‘.dvi’) file. This contains commands that can be translated into commands for a variety of output devices. You can view such ‘.dvi’ output of L^{A}TEX by using a program such as xdvi (display directly) or dvips (convert to PostScript).

2. If invoked as pdflatex, a “Portable Document Format” (‘.pdf’) file. Typically, this is a self-contained file, with all fonts and images embedded. This can be very useful, but it does make the output much larger than the ‘.dvi’ produced from the same document.

There are other less-common variants of L^{A}TEX (and TEX) as well, which can produce HTML, XML, and other things.

2. A “transcript” or ‘.log’ file that contains summary information and diagnostic messages for any errors discovered in the input file.

3. An “auxiliary” or ‘.aux’ file. This is used by L^{A}TEX itself, for things such as sectioning.

A L^{A}TEX command begins with the command name, which consists of a \ followed by either (a) a string of letters or (b) a single non-letter. Arguments contained in square brackets, [], are optional while arguments contained in braces, {}, are required.

L^{A}TEX is case sensitive. Enter all commands in lower case unless explicitly directed to do otherwise.

2 Starting & ending

A minimal input file looks like the following:

     \documentclass{class}
     \begin{document}
     your text
     \end{document}

where the class is a valid document class for L^{A}TEX. See Chapter 3 [Document classes], page 4, for details of the various document classes available locally.

You may include other L^{A}TEX commands between the \documentclass and the \begin{document} commands (this area is called the preamble).

3 Document classes

The class of a given document is defined with the command:

     \documentclass[options]{class}

The \documentclass command must be the first command in a L^{A}TEX source file.

Built-in L^{A}TEX document class names are (many other document classes are available as add-ons; see Chapter 1 [Overview], page 2):

     article  report  book  letter  slides

Standard options are described below.

3.1 Document class options

You can specify so-called global options or class options to the \documentclass command by enclosing them in square brackets as usual. To specify more than one option, separate them with a comma:

     \documentclass[option1,option2,...]{class}

Here is the list of the standard class options.

All of the standard classes except slides accept the following options for selecting the typeface size (default is 10pt):

     10pt  11pt  12pt

All of the standard classes accept these options for selecting the paper size (default is

letterpaper):
     a4paper a5paper b5paper executivepaper legalpaper letterpaper

Miscellaneous other options:

draft, final

mark/do not mark overfull boxes with a big black box; default is final.

fleqn Put displayed formulas flush left; default is centered.

landscape

Selects landscape format; default is portrait.

leqno Put equation numbers on the left side of equations; default is the right side. openbib Use “open” bibliography format.

titlepage, notitlepage

Specifies whether the title page is separate; default depends on the class. These options are not available with the slides class:

onecolumn
twocolumn

Typeset in one or two columns; default is onecolumn.

oneside

twoside Selects one- or two-sided layout; default is oneside, except for the book class. The \evensidemargin (\oddsidemargin parameter determines the distance on even (odd) numbered pages between the left side of the page and the text’s left margin. The defaults vary with the paper size and whether oneor two-side layout is selected. For one-sided printing the text is centered, for two-sided, \oddsidemargin is 40% of the difference between \paperwidth and

          \textwidth with \evensidemargin the remainder.
openright

openany Determines if a chapter should start on a right-hand page; default is openright for book.

The slides class offers the option clock for printing the time at the bottom of each note.

Additional packages are loaded like this:

     \usepackage[options]{pkg}

To specify more than one pkg, you can separate them with a comma, or use multiple

\usepackage commands.

Any options given in the \documentclass command that are unknown by the selected document class are passed on to the packages loaded with \usepackage.

4 Typefaces

Two important aspects of selecting a font are specifying a size and a style. The L^{A}TEX commands for doing this are described here.

4.1 Font styles

The following type style commands are supported by L^{A}TEX.

These commands are used like \textit{italic text}. The corresponding command in parenthesis is the “declaration form”, which takes no arguments. The scope of the declaration form lasts until the next type style command or the end of the current group.

The declaration forms are cumulative; i.e., you can say either \sffamily\bfseries or \bfseries\sffamily to get bold sans serif.

You can also use the environment form of the declaration forms; for instance,

\begin{ttfamily}...\end{ttfamily}.
\textrm (\rmfamily)

Roman.

\textit (\itshape)

Italics.

\emph Emphasis (switches between \textit and \textrm).

\textmd (\mdseries)

Medium weight (default).

\textbf (\bfseries)

Boldface.

\textup (\upshape)

Upright (default). The opposite of slanted.

\textsl (\slshape)

Slanted.

\textsf (\sffamily)

Sans serif.

\textsc (\scshape)

Small caps.

\texttt (\ttfamily)

Typewriter.

\textnormal (\normalfont)

Main document font.

\mathrm Roman, for use in math mode.

\mathbf Boldface, for use in math mode.

\mathsf Sans serif, for use in math mode.

\mathtt Typewriter, for use in math mode.

\mathit

(\mit) Italics, for use in math mode.

\mathnormal

For use in math mode, e.g. inside another type style declaration.

\mathcal ‘Calligraphic’ letters, for use in math mode.

In addition, the command \mathversion{bold} can be used for switching to bold letters and symbols in formulas. \mathversion{normal} restores the default.

L^{A}TEX also provides these commands, which unconditionally switch to the given style, that is, are not cumulative. They are used differently than the above commands, too: {\cmd ...} instead of \cmd{...}. These are two very different things. \bf Switch to bold face. \cal Switch to calligraphic letters for math. \em Emphasis (italics within roman, roman within italics). \it Italics. \rm Roman. \sc Small caps. \sf Sans serif. \sl Slanted (oblique). \tt Typewriter (monospace, fixed-width).

4.2 Font sizes

The following standard type size commands are supported by L^{A}TEX. The table shows the command name and the corresponding actual font size used (in points) with the ‘10pt’, ‘11pt’, and ‘12pt’ document size options, respectively (see Section 3.1 [Document class options], page 4).

Command  10pt   11pt   12pt
\tiny  5  6  6
\scriptsize  7  8  8
\footnotesize  8  9  10

\small  9  10  10.95
\normalsize (default) 10  10.95 12
\large 12 12 14.4
\Large  14.4 14.4 17.28
\LARGE  17.28 17.28 20.74
\huge  20.74 20.74 24.88
\Huge  24.88 24.88 24.88

The commands as listed here are “declaration forms”. The scope of the declaration form lasts until the next type style command or the end of the current group. You can also use the environment form of these commands; for instance, \begin{tiny}...\end{tiny}.

4.3 Low-level font commands

These commands are primarily intended for writers of macros and packages. The commands listed here are only a subset of the available ones.

\fontencoding{enc}

Select font encoding. Valid encodings include OT1 and T1.

\fontfamily{family}

Select font family. Valid families include:

• cmr for Computer Modern Roman

• cmss for Computer Modern Sans Serif

• cmtt for Computer Modern Typewriter and numerous others.

\fontseries{series}

Select font series. Valid series include:

• m Medium (normal)

• b Bold

• c Condensed

• bc Bold condensed

• bx Bold extended and various other combinations.

\fontshape{shape}

Select font shape. Valid shapes are:

• n Upright (normal)

• it Italic

• sl Slanted (oblique)

• sc Small caps

• ui Upright italics

• ol Outline The two last shapes are not available for most font families.

\fontsize{size}{skip}

Set font size. The first parameter is the font size to switch to and the second is the line spacing to use; this is stored in a parameter named \baselineskip. The unit of both parameters defaults to pt. The default \baselineskip for the Computer Modern typeface is 1.2 times the \fontsize. The line spacing is also multiplied by the value of the \baselinestretch parameter when the type size changes; the default is 1. However, the best way to “double space” a document, if you should be unlucky enough to have to produce such, is to use the setspace package; see

http://www.tex.ac.uk/cgi-bin/texfaq2html?label=linespace.

\selectfont

The changes made by calling the four font commands described above do not come into effect until \selectfont is called.

\usefont{enc}{family}{series}{shape}
The same as invoking \fontencoding, \fontfamily, \fontseries and

\fontshape with the given parameters, followed by \selectfont.

5 Layout

Miscellaneous commands for controlling the general layout of the page.

5.1 \onecolumn

The \onecolumn declaration starts a new page and produces single-column output. This is the default.

5.2 \twocolumn

Synopsis:

     \twocolumn[text1col]

The \twocolumn declaration starts a new page and produces two-column output. If the optional text1col argument is present, it is typeset in one-column mode before the two-column typesetting starts.

These parameters control typesetting in two-column output:

\columnsep

The distance between columns (35pt by default).

\columnseprule

The width of the rule between columns; the default is 0pt, so there is no rule.

\columnwidth

The width of the current column; this is equal to \textwidth in single-column text.

These parameters control float behavior in two-column output:

\dbltopfraction

Maximum fraction at the top of a two-column page that may be occupied by floats. Default ‘.7’, can be usefully redefined to (say) ‘.9’ to avoid going to float pages so soon.

\dblfloatpagefraction

The minimum fraction of a float page that must be occupied by floats, for a two-column float page. Default ‘.5’.

\dblfloatsep

Distance between floats at the top or bottom of a two-column float page. De-

          fault ‘12pt plus2pt minus2pt’ for ‘10pt’ and ‘11pt’ documents, ‘14pt plus2pt
          minus4pt’ for ‘12pt’.
\dbltextfloatsep

Distance between a multi-column float at the top or bottom of a page and the

          main text. Default ‘20pt plus2pt minus4pt’.

5.3 \flushbottom

The \flushbottom declaration makes all text pages the same height, adding extra vertical space where necessary to fill out the page.

This is the default if twocolumn mode is selected (see Section 3.1 [Document class options], page 4).

5.4 \raggedbottom

The \raggedbottom declaration makes all pages the natural height of the material on that page. No rubber lengths will be stretched.

5.5 Page layout parameters

\headheight

Height of the box that contains the running head. Default is ‘30pt’, except in the book class, where it varies with the type size.

\headsep Vertical distance between the bottom of the header line and the top of the main text. Default is ‘25pt’, except in the book class, where it varies with the type size.

\footskip

Distance from the baseline of the last line of text to the baseline of the page footer. Default is ‘30pt’, except in the book class, where it varies with the type size.

\linewidth

Width of the current line; the default varies with the font size, paper width, two-column mode, etc. For an article document in ‘10pt’, it’s set to ‘345pt’; in two-column mode, that becomes ‘229.5pt’.

\textheight

The normal vertical height of the page body; the default varies with the font size, document class, etc. For an article or report document in ‘10pt’, it’s

set to ‘43\baselineskip’; for book, it’s ‘41\baselineskip’. For ‘11pt’, it’s
‘38\baselineskip’ and for ‘12pt’, ‘36\baselineskip’.
\textwidth

The normal horizontal width of the page body; the default varies as usual. For an article or report document, it’s ‘345pt’ at ‘10pt’, ‘360pt’ at ‘11pt’, and ‘390pt’ at ‘12pt’. For a book document, it’s ‘4.5in’ at ‘10pt’, and ‘5in’ at

‘11pt’ or ‘12pt’.
\topmargin

Space between the top of the TEX page (one inch from the top of the paper, by default) and the top of the header. The default is computed based on

many other parameters: \paperheight − 2in − \headheight − \headsep −

\textheight − \footskip, and then divided by two.

\topskip Minimum distance between the top of the page body and the baseline of the first line of text. For the standard clases, the default is the same as the font size, e.g., ‘10pt’ at ‘10pt’.

6 Sectioning

Sectioning commands provide the means to structure your text into units:

\part

\chapter (report and book class only)

\section
\subsection
\subsubsection
\paragraph
\subparagraph

All sectioning commands take the same general form, e.g.,

     \chapter[toctitle]{title}

In addition to providing the heading title in the main text, the section title can appear in two other places:

1. The table of contents.

2. The running head at the top of the page.

You may not want the same text in these places as in the main text. To handle this, the sectioning commands have an optional argument toctitle that, when given, specifies the text for these other places.

Also, all sectioning commands have *-forms that print title as usual, but do not include a number and do not make an entry in the table of contents. For instance:

     \section*{Preamble}

The \appendix command changes the way following sectional units are numbered. The \appendix command itself generates no text and does not affect the numbering of parts. The normal use of this command is something like

     \chapter{A Chapter}
     ...
     \appendix
     \chapter{The First Appendix}

7 Cross references

One reason for numbering things like figures and equations is to refer the reader to them, as in “See Figure 3 for more details.”

7.1 \label

Synopsis:

     \label{key}

A \label command appearing in ordinary text assigns to key the number of the current sectional unit; one appearing inside a numbered environment assigns that number to key.

A key name can consist of any sequence of letters, digits, or punctuation characters. Upper and lowercase letters are distinguished.

To avoid accidentally creating two labels with the same name, it is common to use labels consisting of a prefix and a suffix separated by a colon or period. Some conventionally-used prefixes: ch for chapters sec for lower-level sectioning commands fig for figures tab for tables eq for equations

Thus, a label for a figure would look like fig:snark or fig.snark.

7.2 \pageref{key}

Synopsis:

     \pageref{key}

The \pageref{key} command produces the page number of the place in the text where the corresponding \label{key} command appears.

7.3 \ref{key}

Synopsis:

     \ref{key}

The \ref command produces the number of the sectional unit, equation, footnote, figure, …, of the corresponding \label command (see Section 7.1 [\label], page 13). It does not produce any text, such as the word ‘Section’ or ‘Figure’, just the bare number itself.

8 Environments

L^{A}TEX provides many environments for marking off certain text. Each environment begins and ends in the same manner:

     \begin{envname}
     ...
     \end{envname}
8.1 abstract

Synopsis:

     \begin{abstract}
     ...
     \end{abstract}

Environment for producing an abstract, possibly of multiple paragraphs.

8.2 array

Synopsis:

     \begin{array}{template}
     col1 text&col1 text&coln}\\
     ...
     \end{array}

Math arrays are produced with the array environment, normally within an equation environment (see Section 8.9 [equation], page 17). It has a single mandatory template argument describing the number of columns and the alignment within them. Each column col is specified by a single letter that tells how items in that row should be formatted, as follows:

c centered

l flush left

r flush right

Column entries are separated by &. Column entries may include other L^{A}TEX commands. Each row of the array is terminated with \\.

In the template, the construct @{text} puts text between columns in each row.

Here’s an example:

     \begin{equation}
       \begin{array}{lrc}
       left1 & right1 & centered1 \\
       left2 & right2 & centered2 \\
       \end{array}
     \end{equation}

The \arraycolsep parameter defines half the width of the space separating columns; the default is ‘5pt’. See Section 8.23 [tabular], page 28, for other parameters which affect formatting in array environments, namely \arrayrulewidth and \arraystretch.

The array environment can only be used in math mode.

8.3 center

Synopsis:

     \begin{center}
     line1 \\
     line2 \\
     \end{center}

The center environment allows you to create a paragraph consisting of lines that are centered within the left and right margins on the current page. Each line is terminated with the string \\.

8.3.1 \centering

The \centering declaration corresponds to the center environment. This declaration can be used inside an environment such as quote or in a parbox. Thus, the text of a figure or table can be centered on the page by putting a \centering command at the beginning of the figure or table environment.

Unlike the center environment, the \centering command does not start a new paragraph; it simply changes how L^{A}TEX formats paragraph units. To affect a paragraph unit’s format, the scope of the declaration must contain the blank line or \end command (of an environment such as quote) that ends the paragraph unit.

Here’s an example:

     \begin{quote}
     \centering
     first line \\
     second line \\
     \end{quote}
8.4 description

Synopsis:

     \begin{description}
     \item [label1] item1
     \item [label2] item2
     ...
     \end{description}

The description environment is used to make labelled lists. Each label is typeset in bold, flush right. The item text may contain multiple paragraphs.

Another variation: since the bold style is applied to the labels, if you typeset a label in typewriter using \texttt, you’ll get bold typewriter: \item[\texttt{bold and typewriter}]. This may be too bold, among other issues. To get just typewriter, use \tt, which resets all other style variations: \item[{\tt plain typewriter}].

For details about list spacing, see Section 8.13 [itemize], page 19.

8.5 displaymath

Synopsis:

     \begin{displaymath}
     math
     \end{displaymath}

or

     \[math\]

The displaymath environment (\[...\] is a synonym) typesets the math text on its own line, centered by default. The global fleqn option makes equations flush left; see Section 3.1 [Document class options], page 4.

No equation number is added to displaymath text; to get an equation number, use the equation environment (see Section 8.9 [equation], page 17).

8.6 document

The document environment encloses the body of a document. It is required in every L^{A}TEX document. See Chapter 2 [Starting & ending], page 3.

8.7 enumerate

Synopsis:

     \begin{enumerate}
     \item item1
     \item item2
     ...
     \end{enumerate}

The enumerate environment produces a numbered list. Enumerations can be nested within one another, up to four levels deep. They can also be nested within other paragraph-making environments, such as itemize (see Section 8.13 [itemize], page 19) and description (see Section 8.4 [description], page 15).

Each item of an enumerated list begins with an \item command. There must be at least one \item command within the environment.

By default, the numbering at each level is done like this:

1. 1., 2., …

2. (a), (b), …

3. i., ii., …

4. A., B., …

The enumerate environment uses the counters \enumi through \enumiv counters (see Chapter 11 [Counters], page 38). If the optional argument to \item is given, the counter is not incremented for that item.

The enumerate environment uses the commands \labelenumi through \labelenumiv to produce the default label. So, you can use \renewcommand to change the labels (see Section 10.1 [\newcommand & \renewcommand], page 35). For instance, to have the first level use uppercase letters:

     \renewcommand{\labelenumi}{\Alph{enumi}}

8.8 eqnarray
\begin{eqnarray} (or eqnarray*)
formula1 \\
formula2 \\
...
\end{eqnarray}

The eqnarray environment is used to display a sequence of equations or inequalities. It is very much like a three-column array environment, with consecutive rows separated by \\ and consecutive items within a row separated by an &.

\\* can also be used to separate equations, with its normal meaning of not allowing a page break at that line.

An equation number is placed on every line unless that line has a \nonumber command. Alternatively, The *-form of the environment (\begin{eqnarray*} ... \end{eqnarray*}) will omit equation numbering entirely, while otherwise being the same as eqnarray.

The command \lefteqn is used for splitting long formulas across lines. It typesets its argument in display style flush left in a box of zero width.

8.9 equation

Synopsis:

\begin{equation}
math
\end{equation}

The equation environment starts a displaymath environment (see Section 8.5 [displaymath], page 15), e.g., centering the math text on the page, and also places an equation number in the right margin.

8.10 figure
\begin{figure}[placement]
figbody
\label{label}
\caption[loftitle]{text}
\end{figure}

Figures are objects that are not part of the normal text, and are instead “floated” to a convenient place, such as the top of a page. Figures will not be split between two pages.

The optional argument [placement] determines where L^{A}TEX will try to place your figure. There are four places where L^{A}TEX can possibly put a float:

1. h (Here) - at the position in the text where the figure environment appears.

2. t (Top) - at the top of a text page.

3. b (Bottom) - at the bottom of a text page.

4. p (Page of floats) - on a separate float page, which is a page containing no text, only floats.

The standard report and article classes use the default placement tbp.

The body of the figure is made up of whatever text, L^{A}TEX commands, etc. you wish. The \caption command specifies caption text for the figure. The caption is numbered

by default. If loftitle is present, it is used in the list of figures instead of text (see Section 23.1 [Tables of contents], page 71).

The maximum fraction of the page allowed to be occuped by floats at the bottom; default ‘.3’.

\floatpagefraction

The minimum fraction of a float page that must be occupied by floats; default ‘.5’.

\floatsep

Space between floats at the top or bottom of a page; default ‘12pt plus2pt

minus2pt’.
\intextsep

Space above and below a float in the middle of the main text; default ‘12pt

plus2pt minus2pt’ for ‘10pt’ and ‘11pt’ styles, ‘14pt plus4pt minus4pt’ for
‘12pt’.
\textfloatsep

Space between the last (first) float at the top (bottom) of a page; default ‘20pt

plus2pt minus4pt’.
\textfraction

Minimum fraction of a page that must be text; if floats take up too much space to preserve this much text, floats will be moved to a different page. The default is ‘.2’.

\topfraction

Maximum fraction at the top of a page that may be occupied before floats; default is ‘.7’.

8.11 flushleft
\begin{flushleft}
line1 \\
line2 \\
...
\end{flushleft}

The flushleft environment allows you to create a paragraph consisting of lines that are flush to the left-hand margin and ragged right Each line must be terminated with the string \\.

8.11.1 \raggedright

The \raggedright declaration corresponds to the flushleft environment. This declaration can be used inside an environment such as quote or in a parbox.

Unlike the flushleft environment, the \raggedright command does not start a new paragraph; it only changes how L^{A}TEX formats paragraph units. To affect a paragraph unit’s format, the scope of the declaration must contain the blank line or \end command that ends the paragraph unit.

8.12 flushright
     \begin{flushright}
     line1 \\
     line2 \\
     ...
     \end{flushright}

The flushright environment allows you to create a paragraph consisting of lines that are flush to the right-hand margin and ragged left. Each line must be terminated with the string \\.

8.12.1 \raggedleft

The \raggedleft declaration corresponds to the flushright environment. This declaration can be used inside an environment such as quote or in a parbox.

Unlike the flushright environment, the \raggedleft command does not start a new paragraph; it only changes how L^{A}TEX formats paragraph units. To affect a paragraph unit’s format, the scope of the declaration must contain the blank line or \end command that ends the paragraph unit.

8.13 itemize

Synopsis:

     \begin{itemize}
     \item item1
     \item item2
     ...
     \end{itemize}

The itemize environment produces an “unordered”, “bulleted” list. Itemizations can be nested within one another, up to four levels deep. They can also be nested within other paragraph-making environments, such as enumerate (see Section 8.7 [enumerate], page 16).

Each item of an itemize list begins with an \item command. There must be at least one \item command within the environment.

By default, the marks at each level look like this:

1. •

2. -- (bold en-dash)

3. *

4. · (centered dot)

The itemize environment uses the commands \labelitemi through \labelitemiv to produce the default label. So, you can use \renewcommand to change the labels. For instance, to have the first level use diamonds:

     \renewcommand{\labelitemi}{$\diamond$}

The \leftmargini through \leftmarginvi parameters define the distance between the left margin of the enclosing environment and the left margin of the list. By convention, \leftmargin is set to the appropriate \leftmarginN when a new level of nesting is entered.

The defaults vary from ‘.5em’ (highest levels of nesting) to ‘2.5em’ (first level), and are a bit reduced in two-column mode. This example greatly reduces the margin space for outermost lists:

     \setlength{\leftmargini}{1.25em} % default 2.5em

Some parameters that affect list formatting:

\itemindent

Extra indentation before each item in a list; default zero.

\labelsep

Space between the label and text of an item; default ‘.5em’.

\labelwidth

Width of the label; default ‘2em’, or ‘1.5em’ in two-column mode.

\listparindent

Extra indentation added to second and subsequent paragraphs within a list item; default ‘0pt’.

\rightmargin

Horizontal distance between the right margin of the list and the enclosing environment; default ‘0pt’, except in the quote, quotation, and verse environments, where it is set equal to \leftmargin.

Parameters affecting vertical spacing between list items (rather loose, by default).

\itemsep Vertical space between items. The default is 2pt plus1pt minus1pt for 10pt

          documents, 3pt plus2pt minus1pt for 11pt, and 4.5pt plus2pt minus1pt for
          12pt.

\parsep Extra vertical space between paragraphs within a list item. Defaults are the

          same as \itemsep.

\topsep Vertical space between the first item and the preceding paragraph. The default

          is 4pt plus2pt minus2pt for 10pt documents, 6pt plus2pt minus2pt for 11pt,
          and 9pt plus3pt minus5pt for 12pt.
\partopsep

Extra space added to \topsep when the list environment starts a paragraph.

          The default is 2pt plus1pt minus1pt for 10pt documents, 3pt plus1pt
          minus1pt for 11pt, and 3pt plus2pt minus2pt for 12pt.

\topsep Extra vertical space added before an initial and after a final list item. Its value is changed with list level and font size changes; for instance, within a first-level

          list at ‘10pt’, it is ‘4pt plus2pt minus2pt’.

Especially for lists with short items, it may be desirable to elide space between items. Here is an example defining an itemize* environment with no extra spacing between items, or between paragraphs within a single item (\parskip is not list-specific, see Section 15.3 [\parskip], page 44):

     \newenvironment{itemize*}%
       {\begin{itemize}%
         \setlength{\itemsep}{0pt}%
         \setlength{\parsep}{0pt}}%
         \setlength{\parskip}{0pt}}%
       {\end{itemize}}

8.14 letter environment: writing letters

This environment is used for creating letters. See Chapter 24 [Letters], page 73.

8.15 list

The list environment is a generic environment which is used for defining many of the more specific environments. It is seldom used in documents, but often in macros.

     \begin{list}{labeling}{spacing}
     \item item1
     \item item2
     ...
     \end{list}

The mandatory labeling argument specifies how items should be labelled (unless the optional argument is supplied to \item). This argument is a piece of text that is inserted in a box to form the label. It can and usually does contain other L^{A}TEX commands.

The mandatory spacing argument contains commands to change the spacing parameters for the list. This argument will most often be empty, i.e., {}, which leaves the default spacing.

8.16 math

Synopsis:

     \begin{math}
     math
     \end{math}

The math environment inserts the given math within the running text. \(...\)) and $...$ are synonyms. See Chapter 16 [Math formulas], page 46.

8.17 minipage
     \begin{minipage}[position]{width}
     text
     \end{minipage}

The minipage environment typesets its body text in a block that will not be broken across pages (similar to the \parbox command, see Section 20.5 [\parbox], page 65). You may use other paragraph-making environments inside a minipage (unlike \parbox).

It takes an optional position argument for placing text, and a mandatory width argument for specifying the line width.

By default, paragraphs are not indented in the minipage environment. You can restore indentation with a command such as \setlength{\parindent}{1pc} command.

Footnotes in a minipage environment are handled in a way that is particularly useful for putting footnotes in figures or tables. A \footnote or \footnotetext command puts the footnote at the bottom of the minipage instead of at the bottom of the page, and it uses the \mpfootnote counter instead of the ordinary footnote counter (see Chapter 11 [Counters], page 38).

However, don’t put one minipage inside another if you are using footnotes; they may wind up at the bottom of the wrong minipage.

8.18 picture
     \begin{picture}(width,height)(x offset,y offset)
     ... picture commands ...
     \end{picture}

The picture environment allows you to create just about any kind of picture you want containing text, lines, arrows and circles. You tell L^{A}TEX where to put things in the picture by specifying their coordinates. A coordinate is a number that may have a decimal point and a minus sign—a number like 5, 0.3 or -3.1416. A coordinate specifies a length in multiples of the unit length \unitlength, so if \unitlength has been set to 1cm, then the coordinate 2.54 specifies a length of 2.54 centimeters. You should only change the value of \unitlength, using the \setlength command, outside of a picture environment.

A position is a pair of coordinates, such as (2.4,-5), specifying the point with xcoordinate 2.4 and y-coordinate -5. Coordinates are specified in the usual way with respect to an origin, which is normally at the lower-left corner of the picture. Note that when a position appears as an argument, it is not enclosed in braces; the parentheses serve to delimit the argument.

The picture environment has one mandatory argument, which is a position. It specifies the size of the picture. The environment produces a rectangular box with width and height determined by this argument’s x- and y-coordinates.

The picture environment also has an optional position argument, following the size argument, that can change the origin. (Unlike ordinary optional arguments, this argument is not contained in square brackets.) The optional argument gives the coordinates of the point at the lower-left corner of the picture (thereby determining the origin). For example, if \unitlength has been set to 1mm, the command

        \begin{picture}(100,200)(10,20)

produces a picture of width 100 millimeters and height 200 millimeters, whose lower-left corner is the point (10,20) and whose upper-right corner is therefore the point (110,220). When you first draw a picture, you will omit the optional argument, leaving the origin at the lower-left corner. If you then want to modify your picture by shifting everything, you just add the appropriate optional argument.

The environment’s mandatory argument determines the nominal size of the picture. This need bear no relation to how large the picture really is; L^{A}TEX will happily allow you to put things outside the picture, or even off the page. The picture’s nominal size is used by L^{A}TEX in determining how much room to leave for it.

Everything that appears in a picture is drawn by the \put command. The command

     \put (11.3,-.3){...}

puts the object specified by ... in the picture, with its reference point at coordinates (11.3, −.3). The reference points for various objects will be described below.

The \put command creates an LR box. You can put anything that can go in an \mbox (see Section 20.1 [\mbox], page 64) in the text argument of the \put command. When you do this, the reference point will be the lower left corner of the box.

The picture commands are described in the following sections.

8.18.1 \circle
     \circle[*]{diameter}

The \circle command produces a circle with a diameter as close to the specified one as possible. The *-form of the command draws a solid circle.

Circles up to 40 pt can be drawn.

8.18.2 \makebox
\makebox(width,height)[position]{...}

The \makebox command for the picture environment is similar to the normal \makebox command except that you must specify a width and height in multiples of \unitlength.

The optional argument, [position], specifies the quadrant that your text appears in. You may select up to two of the following:

• t - Moves the item to the top of the rectangle

• b - Moves the item to the bottom

• l - Moves the item to the left

• r - Moves the item to the right

See Section 20.4 [\makebox], page 64.

8.18.3 \framebox

Synopsis:

     \framebox(width,height)[pos]{...}

The \framebox command is like \makebox (see previous section), except that it puts a frame around the outside of the box that it creates.

The \framebox command produces a rule of thickness \fboxrule, and leaves a space \fboxsep between the rule and the contents of the box.

8.18.4 \dashbox

Draws a box with a dashed line. Synopsis:

     \dashbox{dlen}(rwidth,rheight)[pos]{text}

\dashbox creates a dashed rectangle around text in a picture environment. Dashes are dlen units long, and the rectangle has overall width rwidth and height rheight. The text is positioned at optional pos.

A dashed box looks best when the rwidth and rheight are multiples of the dlen.

8.18.5 \frame

Synopsis:

     \frame{text}

The \frame command puts a rectangular frame around text. The reference point is the bottom left corner of the frame. No extra space is put between the frame and the object.

8.18.6 \line

Synopsis:

     \line(xslope,yslope){length}

The \line command draws a line with the given length and slope xslope/yslope.

Standard L^{A}TEX can only draw lines with slope = x/y, where x and y have integer values from −6 through 6. For lines of any slope, not to mention other shapes, see the curve2e and many many other packages on CTAN.

8.18.7 \linethickness

The \linethickness{dim} command declares the thickness of horizontal and vertical lines in a picture environment to be dim, which must be a positive length.

\linethickness does not affect the thickness of slanted lines, circles, or the quarter circles drawn by \oval.

8.18.8 \thicklines

The \thicklines command is an alternate line thickness for horizontal and vertical lines in a picture environment; cf. Section 8.18.7 [\linethickness], page 24 and Section 8.18.9 [\thinlines], page 24.

8.18.9 \thinlines

The \thinlines command is the default line thickness for horizontal and vertical lines in a picture environment; cf. Section 8.18.7 [\linethickness], page 24 and Section 8.18.8 [\thicklines], page 24.

8.18.10 \multiput

Synopsis:

     \multiput(x,y)(delta_x,delta_y){n}{obj}

The \multiput command copies the object obj in a regular pattern across a picture. obj is first placed at position (x, y), then at (x + δx, y + δy), and so on, n times.

8.18.11 \oval

Synopsis:

     \oval(width,height)[portion]

The \oval command produces a rectangle with rounded corners. The optional argument portion allows you to select part of the oval via the following:

t selects the top portion;

b selects the bottom portion; r selects the right portion;

l selects the left portion.

The “corners” of the oval are made with quarter circles with a maximum radius of 20 pt, so large “ovals” will look more like boxes with rounded corners.

8.18.12 \put

\put(x coord,y coord){ ... }

The \put command places the item specified by the mandatory argument at the given coordinates.

8.18.13 \shortstack

Synopsis:

     \shortstack[position]{...\\...\\...}

The \shortstack command produces a stack of objects. The valid positions are:

r Move the objects to the right of the stack.

l Move the objects to the left of the stack

c Move the objects to the centre of the stack (default)

Objects are separated with \\.

8.18.14 \vector

Synopsis:

     \vector(x-slope,y-slope){length}

The \vector command draws a line with an arrow of the specified length and slope. The x and y values must lie between −4 and +4, inclusive.

8.19 quotation

Synopsis:

     \begin{quotation}
     text
     \end{quotation}

The margins of the quotation environment are indented on both the left and the right. The text is justified at both margins. Leaving a blank line between text produces a new paragraph.

Unlike the quote environment, each paragraph is indented normally.

8.20 quote

Snyopsis:

     \begin{quote}
     text
     \end{quote}

The margins of the quote environment are indented on both the left and the right. The text is justified at both margins. Leaving a blank line between text produces a new paragraph.

Unlike the quotation environment, paragraphs are not indented.

8.21 tabbing

Synopsis:

\begin{tabbing}
row1col1 \= row1col2 \= row1col3 \= row1col4 \\
row2col1 \>                \> row2col3 \\
...
\end{tabbing}

The tabbing environment provides a way to align text in columns. It works by setting tab stops and tabbing to them much as was done on an ordinary typewriter. It is best suited for cases where the width of each column is constant and known in advance.

This environment can be broken across pages, unlike the tabular environment.

The following commands can be used inside a tabbing enviroment:

\\ tabbing End a line.

\= (tabbing)

Sets a tab stop at the current position.

\> (tabbing)

Advances to the next tab stop.

\< Put following text to the left of the local margin (without changing the margin). Can only be used at the start of the line.

\+ Moves the left margin of the next and all the following commands one tab stop to the right, beginning tabbed line if necessary.

\- Moves the left margin of the next and all the following commands one tab stop to the left, beginning tabbed line if necessary.

\' (tabbing)

Moves everything that you have typed so far in the current column, i.e. everything from the most recent \>, \<, \', \\, or \kill command, to the right of the previous column, flush against the current column’s tab stop.

\` (tabbing)

Allows you to put text flush right against any tab stop, including tab stop 0. However, it can’t move text to the right of the last column because there’s no tab stop there. The \` command moves all the text that follows it, up to the \\ or \end{tabbing} command that ends the line, to the right margin of the tabbing environment. There must be no \> or \' command between the \` and the command that ends the line.

\a (tabbing)

In a tabbing environment, the commands \=, \' and \` do not produce accents as usual (see Section 21.3 [Accents], page 68). Instead, the commands \a=, \a' and \a` are used.

\kill Sets tab stops without producing text. Works just like \\ except that it throws away the current line instead of producing output for it. The effect of any \=, \+ or \- commands in that line remain in effect.

\poptabs Restores the tab stop positions saved by the last \pushtabs.

\pushtabs

Saves all current tab stop positions. Useful for temporarily changing tab stop positions in the middle of a tabbing environment.

\tabbingsep

Distance to left of tab stop moved by \'.

This example typesets a Pascal function in a traditional format:

     \begin{tabbing}
     function \= fact(n : integer) : integer;\\
              \> begin \= \+ \\
                    \> if \= n $>$ 1 then \+ \\
                             fact := n * fact(n-1) \- \\
                       else \+ \\
                             fact := 1; \-\- \\
                 end;\\
     \end{tabbing}
8.22 table

Synopsis:

      \begin{table}[placement]
       body of the table
      \caption{table title}
      \end{table}

Tables are objects that are not part of the normal text, and are usually “floated” to a convenient place, like the top of a page. Tables will not be split between two pages.

The optional argument [placement] determines where L^{A}TEX will try to place your table. There are four places where L^{A}TEX can possibly put a float:

• h: Here - at the position in the text where the table environment appears.

• t: Top - at the top of a text page.

• b: Bottom - at the bottom of a text page.

• p: Page of floats - on a separate float page, which is a page containing no text, only floats.

The standard report and article classes use the default placement [tbp].

The body of the table is made up of whatever text, L^{A}TEX commands, etc., you wish. The \caption command allows you to title your table.

8.23 tabular

Synopsis:

     \begin{tabular}[pos]{cols}
     column 1 entry & column 2 entry ... & column n entry \\
     ...
     \end{tabular}

or

     \begin{tabular*}{width}[pos]{cols}
     column 1 entry & column 2 entry ... & column n entry \\
     ...
     \end{tabular*}

These environments produce a box consisting of a sequence of rows of items, aligned vertically in columns.

\\ must be used to specify the end of each row of the table, except for the last, where it is optional—unless an \hline command (to put a rule below the table) follows.

The mandatory and optional arguments consist of:

width Specifies the width of the tabular* environment. There must be rubber space between columns that can stretch to fill out the specified width.

pos Specifies the vertical position; default is alignment on the centre of the envi-

ronment.

• t - align on top row

• b - align on bottom row

cols Specifies the column formatting. It consists of a sequence of the following

specifiers, corresponding to the sequence of columns and intercolumn material. • l - A column of left-aligned items.

• r - A column of right-aligned items.

• c - A column of centred items.

• | - A vertical line the full height and depth of the environment.

• @{text} - This inserts text in every row. An @-expression suppresses the intercolumn space normally inserted between columns; any desired space between the inserted text and the adjacent items must be included in text. An \extracolsep{wd} command in an @-expression causes an extra space of width wd to appear to the left of all subsequent columns, until countermanded by another \extracolsep command. Unlike ordinary intercolumn space, this extra space is not suppressed by an @-expression. An \extracolsep command can be used only in an @-expression in the cols argument.

• p{wd} - Produces a column with each item typeset in a parbox of width wd, as if it were the argument of a \parbox[t]{wd} command. However, a \\ may not appear in the item, except in the following situations:

1. inside an environment like minipage, array, or tabular.

2. inside an explicit \parbox.

               3. in the scope of a \centering, \raggedright, or \raggedleft decla-

ration. The latter declarations must appear inside braces or an environment when used in a p-column element.

• *{num}{cols} - Equivalent to num copies of cols, where num is any positive integer and cols is any list of column-specifiers, which may contain another

              *-expression.

Parameters that control formatting:

\arrayrulewidth

Thickness of the rule created by |, \hline, and \vline in the tabular and array environments; the default is ‘.4pt’.

\arraystretch

Scaling of spacing between rows in the tabular and array environments; default is ‘1’,for no scaling.

\doublerulesep

Horizontal distance between the vertical rules produced by || in the tabular and array environments; default is ‘2pt’.

\tabcolsep

Half the width of the space between columns; default is ‘6pt’.

These commands can be used inside a tabular environment:

8.23.1 \multicolumn

Synopsis:

     \multicolumn{cols}{pos}{text}

The \multicolumn command makes an entry that spans several columns. The first mandatory argument, cols, specifies the number of columns to span. The second mandatory argument, pos, specifies the formatting of the entry; c for centered, l for flushleft, r for flushright. The third mandatory argument, text, specifies what text to put in the entry.

Here’s an example showing two columns separated by an en-dash; \multicolumn is used for the heading:

     \begin{tabular}{r@{--}l}
     \multicolumn{2}{c}{\bf Unicode}\cr
        0x80&0x7FF   \cr
       0x800&0xFFFF  \cr
     0x10000&0x1FFFF \cr
     \end{tabular}

8.23.2 \cline

Synopsis:

     \cline{i-j}

The \cline command draws horizontal lines across the columns specified, beginning in column i and ending in column j, which are specified in the mandatory argument.

8.23.3 \hline

The \hline command draws a horizontal line the width of the enclosing tabular or array environment. It’s most commonly used to draw a line at the top, bottom, and between the rows of a table.

8.23.4 \vline

The \vline command will draw a vertical line extending the full height and depth of its row. An \hfill command can be used to move the line to the edge of the column. It can also be used in an @-expression.

8.24 thebibliography

Synopsis:

     \begin{thebibliography}{widest-label}
     \bibitem[label]{cite_key}
     ...
     \end{thebibliography}

The thebibliography environment produces a bibliography or reference list.

In the article class, this reference list is labelled “References”; in the report class, it is labelled “Bibliography”. You can change the label (in the standard classes) by redefining the command \refname. For instance, this eliminates it entirely:

     \renewcommand{\refname}{}

The mandatory widest-label argument is text that, when typeset, is as wide as the widest item label produced by the \bibitem commands. It is typically given as 9 for bibliographies with less than 10 references, 99 for ones with less than 100, etc.

8.24.1 \bibitem

Synopsis:

     \bibitem[label]{cite_key}

The \bibitem command generates an entry labelled by label. If the label argument is missing, a number is automatically generated using the enumi counter. The cite key is any sequence of letters, numbers, and punctuation symbols not containing a comma.

This command writes an entry to the ‘.aux’ file containing the item’s cite key and label. When the ‘.aux’ file is read by the \begin{document} command, the item’s label is associated with cite_key, causing references to cite key with a \cite command (see next section) to produce the associated label.

8.24.2 \cite

Synopsis:

     \cite[subcite]{keys

The keys argument is a list of one or more citation keys, separated by commas. This command generates an in-text citation to the references associated with keys by entries in the ‘.aux’ file.

The text of the optional subcite argument appears after the citation. For example, \cite[p.~314]{knuth} might produce ‘[Knuth, p. 314]’.

8.24.3 \nocite
\nocite{key_list}

The \nocite command produces no text, but writes key_list, which is a list of one or more citation keys, on the ‘.aux’ file.

8.24.4 Using BibTEX

If you use the BibTEX program by Oren Patashnik (highly recommended if you need a bibliography of more than a couple of titles) to maintain your bibliography, you don’t use the thebibliography environment (see Section 8.24 [thebibliography], page 30). Instead, you include the lines

     \bibliographystyle{bibstyle}
     \bibliography{bibfile1,bibfile2}

The \bibliographystyle command does not produce any output of its own. Rather, it defines the style in which the bibliography will be produced: bibstyle refers to a file bibstyle‘.bst’, which defines how your citations will look. The standard style names distributed with BibTEX are:

alpha Sorted alphabetically. Labels are formed from name of author and year of publication.

plain Sorted alphabetically. Labels are numeric.

unsrt Like plain, but entries are in order of citation.

abbrv Like plain, but more compact labels.

In addition, numerous other BibTEX style files exist tailored to the demands of various

publications. See http://www.ctan.org/tex-archive/biblio/bibtex/contrib.

The \bibliography command is what actually produces the bibliography. The argument to \bibliography refers to files named ‘bibfile.bib’, which should contain your database in BibTEX format. Only the entries referred to via \cite and \nocite will be listed in the bibliography.

8.25 theorem

Synopsis:

     \begin{theorem}
     theorem-text
     \end{theorem}

The theorem environment produces “Theorem n” in boldface followed by theorem-text, where the numbering possibilities for n are described under \newtheorem (see Section 10.6 [\newtheorem], page 36).

8.26 titlepage

Synopsis:

     \begin{titlepage}
     text
     \end{titlepage}

The titlepage environment creates a title page, i.e., a page with no printed page number or heading. It also causes the following page to be numbered page one. Formatting the title page is left to you. The \today command may be useful on title pages (see Section 21.6 [\today], page 69).

You can use the \maketitle command (see Section 18.1 [\maketitle], page 59) to produce a standard title page without a titlepage environment.

8.27 verbatim

Synopsis:

     \begin{verbatim}
     literal-text
     \end{verbatim}

The verbatim environment is a paragraph-making environment in which L^{A}TEX produces exactly what you type in; for instance the \ character produces a printed ‘\’. It turns L^{A}TEX into a typewriter with carriage returns and blanks having the same effect that they would on a typewriter.

The verbatim uses a monospaced typewriter-like font (\tt).

8.27.1 \verb

Synopsis:

     \verbcharliteral-textchar
     \verb*charliteral-textchar

The \verb command typesets literal-text as it is input, including special characters and spaces, using the typewriter (\tt) font. No spaces are allowed between \verb or \verb* and the delimiter char, which begins and ends the verbatim text. The delimiter must not appear in literal-text.

The *-form differs only in that spaces are printed with a “visible space” character. (Namely, ␣.)

8.28 verse

Synopsis:

     \begin{verse}
     line1 \\
     line2 \\

     ...
     \end{verse}

The verse environment is designed for poetry, though you may find other uses for it. The margins are indented on the left and the right, paragraphs are not indented, and

the text is not justified. Separate the lines of each stanza with \\, and use one or more blank lines to separate the stanzas.

9 Footnotes

Footnotes can be produced in one of two ways. They can be produced with one command, the \footnote command. They can also be produced with two commands, the

\footnotemark and the \footnotetext commands.
9.1 \footnote

Synopsis:

     \footnote[number]{text}

The \footnote command places the numbered footnote text at the bottom of the current page. The optional argument number changes the default footnote number.

This command can only be used in outer paragraph mode; i.e., you cannot use it in sectioning commands like \chapter, in figures, tables or in a tabular environment. (See following sections.)

9.2 \footnotemark

With no optional argument, the \footnotemark command puts the current footnote number in the text. This command can be used in inner paragraph mode. You give the text of the footnote separately, with the \footnotetext command.

This command can be used to produce several consecutive footnote markers referring to the same footnote with

     \footnotemark[\value{footnote}]

after the first \footnote command.

9.3 \footnotetext

Synopsis:

     \footnotetext[number]{text}

The \footnotetext command places text at the bottom of the page as a footnote. This command can come anywhere after the \footnotemark command. The \footnotetext command must appear in outer paragraph mode.

The optional argument number changes the default footnote number.

9.4 Footnote parameters

\footnoterule

Produces the rule separating the main text on a page from the page’s footnotes. Default dimensions: 0.4pt thick (or wide), and 0.4\columnwidth long in the standard document classes (except slides, where it does not appear).

\footnotesep

The height of the strut placed at the beginning of the footnote. By default, this is set to the normal strut for \footnotesize fonts (see Section 4.2 [Font sizes], page 7), therefore there is no extra space between footnotes. This is ‘6.65pt’ for ‘10pt’, ‘7.7pt’ for ‘11pt’, and ‘8.4pt’ for ‘12pt’.

10 Definitions

L^{A}TEX has support for making new commands of many different kinds.

10.1 \newcommand & \renewcommand

\newcommand and \renewcommand define and redefine a command, respectively. Synopses:

       \newcommand{cmd}[nargs]{defn}
     \renewcommand{cmd}[nargs]{defn}
       \newcommand{cmd}[nargs][default]{defn}
     \renewcommand{cmd}[nargs][default]{defn}

cmd The command name beginning with \. For \newcommand, it must not be already defined and must not begin with \end; for \renewcommand, it must already be defined.

nargs An optional integer from 1 to 9 specifying the number of arguments that the command will take. The default is for the command to have no arguments.

default If this optional parameter is present, it means that the command’s first argument is optional. When the new command is called, the default value of the optional argument (i.e., if it is not specified in the call) is the string ‘def’.

defn The text to be substituted for every occurrence of cmd; a construct of the form #n in defn is replaced by the text of the nth argument.

10.2 \newcounter

Synopsis:

     \newcounter{cnt}[countername]

The \newcounter command defines a new counter named cnt. The new counter is initialized to zero.

Given the optional argument [countername], cnt will be reset whenever countername is incremented.

See Chapter 11 [Counters], page 38, for more information about counters.

10.3 \newlength

Synopsis:

     \newlength{\arg}

The \newlength command defines the mandatory argument as a length command with a value of 0in. The argument must be a control sequence, as in \newlength{\foo}. An error occurs if \foo is already defined.

See Chapter 12 [Lengths], page 40, for how to set the new length to a nonzero value, and for more information about lengths in general.

10.4 \newsavebox

Synopsis:

     \newsavebox{cmd}

Defines \cmd, which must be a command name not already defined, to refer to a new bin for storing boxes.

10.5 \newenvironment & \renewenvironment

Synopses:

       \newenvironment{env}[nargs]{begdef}{enddef}
       \newenvironment{env}[nargs][default]{begdef}{enddef}
     \renewenvironment{env}[nargs]{begdef}{enddef}

These commands define or redefine an environment env, that is, \begin{env} ...

\end{env}.

env The name of the environment. For \newenvironment, env must not be an existing environment, and the command \env must be undefined. For \renewenvironment, env must be the name of an existing environment.

nargs An integer from 1 to 9 denoting the number of arguments of the newly-defined environment. The default is no arguments.

default If this is specified, the first argument is optional, and default gives the default value for that argument.

begdef The text expanded at every occurrence of \begin{env}; a construct of the form #n in begdef is replaced by the text of the nth argument.

enddef The text expanded at every occurrence of \end{env}. It may not contain any argument parameters.

10.6 \newtheorem
     \newtheorem{newenv}{label}[within]
     \newtheorem{newenv}[numbered_like]{label}

This command defines a theorem-like environment. Arguments:

newenv The name of the environment to be defined; must not be the name of an existing environment or otherwise defined.

label The text printed at the beginning of the environment, before the number. For example, ‘Theorem’.

numbered like (Optional.) The name of an already defined theorem-like environment; the new environment will be numbered just like numbered like.

within (Optional.) The name of an already defined counter, a sectional unit. The new theorem counter will be reset at the same time as the within counter.

At most one of numbered like and within can be specified, not both.

10.7 \newfont

Synopsis:

     \newfont{cmd}{fontname}

Defines a control sequence \cmd, which must not already be defined, to make fontname be the current font. The file looked for on the system is named ‘fontname.tfm’.

This is a low-level command for setting up to use an individual font. More commonly, fonts are defined in families through ‘.fd’ files.

10.8 \protect

Footnotes, line breaks, any command that has an optional argument, and many more are so-called fragile commands. When a fragile command is used in certain contexts, called moving arguments, it must be preceded by \protect. In addition, any fragile commands within the arguments must have their own \protect.

Some examples of moving arguments are \caption (see Section 8.10 [figure], page 17), \thanks (see Section 18.1 [\maketitle], page 59), and expressions in tabular and array environments (see Section 8.23 [tabular], page 28).

Commands which are not fragile are called robust. They must not be preceded by

\protect.

See also:

http://www-h.eng.cam.ac.uk/help/tpl/textprocessing/teTeX/latex/latex2e-html/fragile.html
http://www.tex.ac.uk/cgi-bin/texfaq2html?label=protect

11 Counters

Everything L^{A}TEX numbers for you has a counter associated with it. The name of the counter is the same as the name of the environment or command that produces the number, except with no \. (enumi–enumiv are used for the nested enumerate environment.) Below is a list of the counters used in L^{A}TEX’s standard document classes to control numbering.

     part            paragraph       figure          enumi
     chapter         subparagraph    table           enumii
     section         page            footnote        enumiii
     subsection      equation        mpfootnote      enumiv
     subsubsection
11.1 \alph \Alph \arabic \roman \Roman \fnsymbol: Printing

counters

All of these commands take a single counter as an argument, for instance, \alph{enumi}.

\alph prints counter using lowercase letters: ‘a’, ‘b’, …

\Alph uses uppercase letters: ‘A’, ‘B’, …

\arabic uses Arabic numbers: ‘1’, ‘2’, …

\roman uses lowercase roman numerals: ‘i’, ‘ii’, …

\roman uses uppercase roman numerals: ‘I’, ‘II’, …

\fnsymbol

prints the value of counter in a specific sequence of nine symbols (conventionally used for labeling footnotes). The value of counter must be between 1 and 9, inclusive. Here are the symbols: ∗ † ‡ § ¶ ∥ ∗∗ †† ‡‡

11.2 \usecounter{counter}

Synopsis:

     \usecounter{counter}

The \usecounter command is used in the second argument of the list environment to specify counter to be used to number the list items.

11.3 \value{counter}

Synopsis:

     \value{counter}

The \value command produces the value of counter. It can be used anywhere L^{A}TEX expects a number, for example:

     \setcounter{myctr}{3}
     \addtocounter{myctr}{1}
     \hspace{\value{myctr}\parindent}

11.4 \setcounter{counter}{value}

Synopsis:

     \setcounter{\counter}{value}

The \setcounter command sets the value of \counter to the value argument.

11.5 \addtocounter{counter}{value}

The \addtocounter command increments counter by the amount specified by the value argument, which may be negative.

11.6 \refstepcounter{counter}

The \refstepcounter command works in the same way as \stepcounter See Section 11.7 [\stepcounter], page 39, except it also defines the current \ref value to be the result of

\thecounter.
11.7 \stepcounter{counter}

The \stepcounter command adds one to counter and resets all subsidiary counters.

11.8 \day \month \year: Predefined counters

L^{A}TEX defines counters for the day of the month (\day, 1–31), month of the year (\month, 1–12), and year (\year, Common Era). When TEX starts up, they are set to the current values on the system where TEX is running. They are not updated as the job progresses.

The related command \today produces a string representing the current day (see Section 21.6 [\today], page 69).

12 Lengths

A length is a measure of distance. Many L^{A}TEX commands take a length as an argument.

12.1 \setlength{\len}{value}

The \setlength sets the value of \len to the value argument, which can be expressed in any units that L^{A}TEX understands, i.e., inches (in), millimeters (mm), points (pt), big points (bp, etc.

12.2 \addtolength{\len}{amount}

The \addtolength command increments a “length command” \len by the amount specified in the amount argument, which may be negative.

12.3 \settodepth
\settodepth{\gnat}{text}

The \settodepth command sets the value of a length command equal to the depth of the text argument.

12.4 \settoheight
\settoheight{\gnat}{text}

The \settoheight command sets the value of a length command equal to the height of the text argument.

12.5 \settowidth{\len}{text}

The \settowidth command sets the value of the command \len to the width of the text argument.

12.6 Predefined lengths

\width
   \height
   \depth
   \totalheight

These length parameters can be used in the arguments of the box-making commands (see Chapter 20 [Boxes], page 64). They specify the natural width etc. of the text in the box. \totalheight equals \height + \depth. To make a box with the text stretched to double the natural size, e.g., say

   \makebox[2\width]{Get a stretcher}

13 Line breaking

The first thing L^{A}TEX does when processing ordinary text is to translate your input file into a sequence of glyphs and spaces. To produce a printed document, this sequence must be broken into lines (and these lines must be broken into pages).

L^{A}TEX usually does the line (and page) breaking for you, but in some environments, you do the line breaking yourself with the \\ command, and you can always manually force breaks.

13.1 \\[*][morespace]

The \\ command tells L^{A}TEX to start a new line. It has an optional argument, morespace, that specifies how much extra vertical space is to be inserted before the next line. This can be a negative amount.

The \\* command is the same as the ordinary \\ command except that it tells L^{A}TEX not to start a new page after the line.

13.2 \obeycr & \restorecr

The \obeycr command makes a return in the input file (‘^^M’, internally) the same as \\ (followed by \relax). So each new line in the input will also be a new line in the output.

\restorecr restores normal line-breaking behavior.

13.3 \newline

The \newline command breaks the line at the present point, with no stretching of the text before it. It can only be used in paragraph mode.

13.4 \- (discretionary hyphen)

The \- command tells L^{A}TEX that it may hyphenate the word at that point. L^{A}TEX is very good at hyphenating, and it will usually find most of the correct hyphenation points, and almost never use an incorrect one. The \- command is used for the exceptional cases.

When you insert \- commands in a word, the word will only be hyphenated at those points and not at any of the hyphenation points that L^{A}TEX might otherwise have chosen.

13.5 \fussy

The declaration \fussy (which is the default) makes TEX picky about line breaking. This usually avoids too much space between words, at the cost of an occasional overfull box.

This command cancels the effect of a previous \sloppy command (see Section 13.6 [\sloppy], page 41.

13.6 \sloppy

The declaration \sloppy makes TEX less fussy about line breaking. This will avoid overfull boxes, at the cost of loose interword spacing.

Lasts until a \fussy command is issued (see Section 13.5 [\fussy], page 41).

13.7 \hyphenation

Synopsis:

     \hyphenation{word-one word-two}

The \hyphenation command declares allowed hyphenation points with a - character in the given words. The words are separated by spaces. TEX will only hyphenate if the word matches exactly, no inflections are tried. Multiple \hyphenation commands accumulate. Some examples (the default TEX hyphenation patterns misses the hyphenations in these words):

     \hyphenation{ap-pen-dix col-umns data-base data-bases}
13.8 \linebreak & \nolinebreak

Synopses:

     \linebreak[priority]
     \nolinebreak[priority]

By default, the \linebreak (\nolinebreak) command forces (prevents) a line break at the current position. For \linebreak, the spaces in the line are stretched out so that it extends to the right margin as usual.

With the optional argument priority, you can convert the command from a demand to a request. The priority must be a number from 0 to 4. The higher the number, the more insistent the request.

14 Page breaking

L^{A}TEX starts new pages asynchronously, when enough material has accumulated to fill up a page. Usually this happens automatically, but sometimes you may want to influence the breaks.

14.1 \cleardoublepage

The \cleardoublepage command ends the current page and causes all figures and tables that have so far appeared in the input to be printed. In a two-sided printing style, it also makes the next page a right-hand (odd-numbered) page, producing a blank page if necessary.

14.2 \clearpage

The \clearpage command ends the current page and causes all figures and tables that have so far appeared in the input to be printed.

14.3 \newpage

The \newpage command ends the current page, but does not clear floats (see \clearpage above).

14.4 \enlargethispage
\enlargethispage{size}
   \enlargethispage*{size}

Enlarge the \textheight for the current page by the specified amount; e.g.

\enlargethispage{\baselineskip} will allow one additional line.

The starred form tries to squeeze the material together on the page as much as possible. This is normally used together with an explicit \pagebreak.

14.5 \pagebreak & \nopagebreak

Synopses:

     \pagebreak[priority]
     \nopagebreak[priority]

By default, the \pagebreak (\nopagebreak) command forces (prevents) a page break at the current position. For \linebreak, the vertical space on the page is stretched out where possible so that it extends to the normal bottom margin.

With the optional argument priority, you can convert the \pagebreak command from a demand to a request. The number must be a number from 0 to 4. The higher the number, the more insistent the request is.

15 Making paragraphs

A paragraph is ended by one or more completely blank lines—lines not containing even a %. A blank line should not appear where a new paragraph cannot be started, such as in math mode or in the argument of a sectioning command.

15.1 \indent

\indent produces a horizontal space whose width equals the width of the \parindent length, the normal paragraph indentation. It is used to add paragraph indentation where it would otherwise be suppressed.

The default value for \parindent is 1em in two-column mode, otherwise 15pt for 10pt documents, 17pt for 11pt, and 1.5em for 12pt.

15.2 \noindent

When used at the beginning of the paragraph, \noindent suppresses any paragraph indentation. It has no effect when used in the middle of a paragraph.

15.3 \parskip

\parskip is a rubber length defining extra vertical space added before each paragraph. The

default is 0pt plus1pt.

15.4 Marginal notes

Synopsis:

     \marginpar[left]{right}

The \marginpar command creates a note in the margin. The first line of the note will have the same baseline as the line in the text where the \marginpar occurs.

When you only specify the mandatory argument right, the text will be placed

• in the right margin for one-sided layout;

• in the outside margin for two-sided layout;

• in the nearest margin for two-column layout.

The command \reversemarginpar places subsequent marginal notes in the opposite (inside) margin. \normalmarginpar places them in the default position.

When you specify both arguments, left is used for the left margin, and right is used for the right margin.

The first word will normally not be hyphenated; you can enable hyphenation there by beginning the node with \hspace{0pt}.

These parameters affect the formatting of the note:

\marginparpush

Minimum vertical space between notes; default ‘7pt’ for ‘12pt’ documents, ‘5pt’ else.

\marginparsep

Horizontal space between the main text and the note; default ‘11pt’ for ‘10pt’ documents, ‘10pt’ else.

\marginparwidth

Width of the note itself; default for a one-sided ‘10pt’ document is ‘90pt’, ‘83pt’ for ‘11pt’, and ‘68pt’ for ‘12pt’; ‘17pt’ more in each case for a two-sided document. In two column mode, the default is ‘48pt’.

16 Math formulas

There are three environments that put L^{A}TEX in math mode:

math For formulas that appear right in the text.

displaymath

For formulas that appear on their own line.

equation The same as the displaymath environment except that it adds an equation number in the right margin.

The math environment can be used in both paragraph and LR mode, but the displaymath and equation environments can be used only in paragraph mode. The math and displaymath environments are used so often that they have the following short forms:

     \(...\) instead of \begin{math}...\end{math}
     \[...\] instead of \begin{displaymath}...\end{displaymath}

In fact, the math environment is so common that it has an even shorter form:

     $ ... $ instead of \(...\)

The \boldmath command changes math letters and symbols to be in a bold font. It is used outside of math mode. Conversely, the \unboldmath command changes math glyphs to be in a normal font; it too is used outside of math mode.

The \displaystyle declaration forces the size and style of the formula to be that of displaymath, e.g., with limits above and below summations. For example

     $\displaystyle \sum_{n=0}^\infty x_n $

16.1 Subscripts & Superscripts

To get an expression exp to appear as a subscript, you just type _{exp}. To get exp to appear as a superscript, you type ^{exp}. L^{A}TEX handles superscripted superscripts and all of that stuff in the natural way. It even does the right thing when something has both a subscript and a superscript.

16.2 Math symbols

L^{A}TEX provides almost any mathematical symbol you’re likely to need. The commands for generating them can be used only in math mode. For example, if you include $\pi$ in your source, you will get the pi symbol (π) in your output.

\|
\aleph  ℵ
\alpha  α

\amalg ⨿ (binary operation)

\angle    ^{}

\approx ≈ (relation)

\ast ∗ (binary operation)

\asymp ≍ (relation)

\backslash

\ (delimiter)

\beta  β
\bigcap ⋂

\bigcirc ◯ (binary operation)

\bigcup ⋃
\bigodot ⨀
\bigoplus

⨁

\bigotimes

⨂

\bigtriangledown

▽ (binary operation)

\bigtriangleup

△ (binary operation)

\bigsqcup

⨆

\biguplus

⨄

\bigcap ⋁
\bigwedge

⋀

\bot  ⊥

\bowtie ▷◁ (relation) \Box (square open box symbol) \bullet • (binary operation) \cap ∩ (binary operation) \cdot · (binary operation)

\chi  χ

\circ ∘ (binary operation)

\clubsuit

♣

\cong ≅ (relation)

\coprod ∐

\cup ∪ (binary operation) \dagger † (binary operation) \dashv ⊣ (relation) \ddagger † (binary operation)

\Delta  Δ
\delta  δ
\Diamond bigger ⋄

\diamond ⋄ (binary operation)

\diamondsuit

♢

\div ÷ (binary operation) \doteq ≐ (relation)

\downarrow

↓ (delimiter)

\Downarrow

⇓ (delimiter)

\ell  ℓ
\emptyset

∅

\epsilon ϵ

\equiv ≡ (relation)

\eta  η
\exists ∃
\flat  ♭
\forall ∀

\frown ⌢ (relation)

\Gamma  Γ
\gamma  γ
\ge  ≥

\geq ≥ (relation)

\gets  ←

\gg ≫ (relation)

\hbar  h̄
\heartsuit

♡

\hookleftarrow

↩

\hookrightarrow

↪

\iff  ⟺
\Im  ℑ

\in ∈ (relation)

\infty  ∞
\int  ∫
\iota  ι

\Join condensed bowtie symbol (relation)

\kappa  κ
\Lambda Λ
\lambda λ
\land  ∧

\langle ⟨ (delimiter) \lbrace { (delimiter) \lbrack [ (delimiter) \lceil ⌈ (delimiter)

\le  ≤
\leadsto
\Leftarrow

⇐

\leftarrow

←

\leftharpoondown

↽

\leftharpoonup

↼

\Leftrightarrow

⇔

\leftrightarrow

↔

\leq ≤ (relation) \lfloor ⌊ (delimiter)

\lhd (left-pointing arrow head) \ll ≪ (relation)

\lnot  ¬
\longleftarrow

⟵

\longleftrightarrow

⟷

\longmapsto

⟼

\longrightarrow

⟶

\lor  ∨
\mapsto ↦
\mho

\mid | (relation) \models ⊧ (relation) \mp ∓ (binary operation)

\mu  μ
\nabla  ∇
\natural ♮
\ne  ≠
\nearrow ↗
\neg  ¬

\neq ≠ (relation) \ni ∋ (relation) \not Overstrike a following operator with a /, as in ≠.

\notin  ∋
\nu  ν
\nwarrow ↖

\odot ⊙ (binary operation)

\oint  ∮
\Omega  Ω
\omega  ω

\ominus ⊖ (binary operation)

\oplus ⊕ (binary operation) \oslash ⊘ (binary operation) \otimes ⊗ (binary operation)

\owns  ∋
\parallel

∥ (relation)

\partial ∂

\perp ⊥ (relation)

\phi  ϕ
\Pi  Π
\pi  π

\pm ± (binary operation) \prec ≺ (relation) \preceq ⪯ (relation)

\prime  ′
\prod  ∏

\propto ∝ (relation)

\Psi  Ψ
\psi  ψ

\rangle ⟩ (delimiter) \rbrace } (delimiter) \rbrack ] (delimiter) \rceil ⌉ (delimiter)

\Re  ℜ
\rfloor ⌋

\rhd (binary operation)

\rho  ρ
\Rightarrow

⇒

\rightarrow

→

\rightharpoondown

⇁

\rightharpoonup

⇀

\rightleftharpoons

⇀ ↽

\searrow ↘
\setminus

\ (binary operation)

\sharp  ♯
\Sigma  Σ
\sigma  σ

\sim ∼ (relation) \simeq ≃ (relation)

\smallint

∫

\smile ⌣ (relation)

\spadesuit

♠

\sqcap ⊓ (binary operation) \sqcup ⊔ (binary operation)

\sqsubset

(relation)

\sqsubseteq

⊑ (relation)

\sqsupset

(relation)

\sqsupseteq

⊒ (relation)

\star ⋆ (binary operation) \subset ⊂ (relation)

\subseteq

⊆ (relation)

\succ ≻ (relation) \succeq ⪰ (relation)

\sum  ∑

\supset ⊃ (relation)

\supseteq

⊇ (relation)

\surd  √

\swarrow ↙
\tau  τ
\theta  θ

\times × (binary operation)

\to  →
\top  ⊤
\triangle

△

\triangleleft

◁ (binary operation)

\triangleright

▷ (binary operation)

\unlhd left-pointing arrowhead with line under (binary operation) \unrhd right-pointing arrowhead with line under (binary operation) \Uparrow ⇑ (delimiter) \uparrow ↑ (delimiter)

\Updownarrow

⇕ (delimiter)

\updownarrow

↕ (delimiter)

\uplus ⊎ (binary operation)

\Upsilon Υ
\upsilon υ
\varepsilon

ε

\varphi φ
\varpi  ϖ
\varrho ϱ
\varsigma

ς

\vartheta

ϑ

\vdash ⊢ (relation) \vee ∨ (binary operation) \Vert ∥ (delimiter)

\vert | (delimiter) \wedge ∧ (binary operation)

\wp  ℘

\wr ≀ (binary operation)

\Xi  Ξ
\xi  ξ
\zeta  ζ

16.3 Math functions

These commands produce roman function names in math mode with proper spacing.

\arccos arccos
\arcsin arcsin
\arctan arctan
\arg  arg

\bmod Binary modulo operator (x mod y)

\cos  cos
\cosh  cosh
\cot  cos
\coth  cosh
\csc  csc
\deg  deg
\det  deg
\dim  dim
\exp  exp
\gcd  gcd
\hom  hom
\inf  inf
\ker  ker
\lg  lg
\lim  lim
\liminf lim inf
\limsup lim sup
\ln  ln

\log  log
\max  max
\min  min

\pmod parenthesized modulus, as in ( (mod 2)^{n} − 1)

\Pr  Pr
\sec  sec
\sin  sin
\sinh  sinh
\sup  sup
\tan  tan
\tanh  tanh

16.4 Math accents

L^{A}TEX provides a variety of commands for producing accented letters in math. These are different from accents in normal text (see Section 21.3 [Accents], page 68). \acute Math acute accent: x́. \breve Math breve accent: x̆. \check Math háček (check) accent: x̌. \ddot Math dieresis accent: ẍ. \dot Math dot accent: ẋ. \grave Math grave accent: x̀. \hat Math hat (circumflex) accent: ˆ x. \imath Math dotless i. \jmath Math dotless j. \tilde Math tilde accent: x̃. \vec Math vector symbol: x⃗. \widehat Math wide hat accent: x̂ + y. \widehat Math wide tilde accent: x̃ + y.

16.5 Spacing in math mode

In a math environment, L^{A}TEX ignores the spaces you type and puts in the spacing according to the normal rules for mathematics texts. If you want different spacing, L^{A}TEX provides the following commands for use in math mode:

\; A thick space (\frac{5}{18} quad).

\:

\> Both of these produce a medium space (\frac{2}{9} quad).

\, A thin space (\frac{1}{6} quad); not restricted to math mode.

\! A negative thin space (−\frac{1}{6} quad).

16.6 Math Miscellany

\* A “discretionary” multiplication symbol, at which a line break is allowed.

\cdots A horizontal ellipsis with the dots raised to the center of the line. As in: ‘⋯’.

\ddots A diagonal ellipsis: ....

\frac{num}{den}

Produces the fraction num divided by den.

eg. \frac{1}{4}

\left delim1 ... \right delim2

The two delimiters need not match; ‘.’ acts as a null delimiter, producing no output. The delimiters are sized according to the math in between. Example:

          \left( \sum_i=1^10 a_i \right].
\overbrace{text}

Generates a brace over text. For example, \overbrace{x + ⋯ + x}^{k times}.

\overline{text}

Generates a horizontal line over tex. For exampe, \overline{x + y}.

\sqrt[root]{arg}

Produces the representation of the square root of arg. The optional argument root determines what root to produce. For example, the cube root of x+y would be typed as $\sqrt[3]\{x+y\}$. In TEX, the result looks like this: ^{3} \sqrt{x} + y.

\stackrel{text}{relation}

Puts text above relation. For example, \stackrel{f}{\longrightarrow}. In TEX, the result looks like this: ⟶^{f}.

\underbrace{math}

Generates math with a brace underneath. In TEX, the result looks like this:

\underbrace{x + y + z}_{> 0}.

\underline{text}

Causes text, which may be either math mode or not, to be underlined. In TEX, the result looks like this: \underline{z} Chapter 16: Math formulas 57 \vdots Produces a vertical ellipsis. In TEX, the result looks like this: ⋮.

17 Modes

When L^{A}TEX is processing your input text, it is always in one of three modes: • Paragraph mode

• Math mode

• Left-to-right mode, called LR mode for short

L^{A}TEX changes mode only when it goes up or down a staircase to a different level, though not all level changes produce mode changes. Mode changes occur only when entering or leaving an environment, or when L^{A}TEX is processing the argument of certain text-producing commands.

“Paragraph mode” is the most common; it’s the one L^{A}TEX is in when processing ordinary text. In that mode, L^{A}TEX breaks your text into lines and breaks the lines into pages. L^{A}TEX is in “math mode” when it’s generating a mathematical formula. In “LR mode”, as in paragraph mode, L^{A}TEX considers the output that it produces to be a string of words with spaces between them. However, unlike paragraph mode, L^{A}TEX keeps going from left to right; it never starts a new line in LR mode. Even if you put a hundred words into an \mbox, L^{A}TEX would keep typesetting them from left to right inside a single box, and then complain because the resulting box was too wide to fit on the line.

L^{A}TEX is in LR mode when it starts making a box with an \mbox command. You can get it to enter a different mode inside the box - for example, you can make it enter math mode to put a formula in the box. There are also several text-producing commands and environments for making a box that put L^{A}TEX in paragraph mode. The box make by one of these commands or environments will be called a parbox. When L^{A}TEX is in paragraph mode while making a box, it is said to be in “inner paragraph mode”. Its normal paragraph mode, which it starts out in, is called “outer paragraph mode”.

18 Page Styles

The \documentclass command determines the size and position of the page’s head and foot. The page style determines what goes in them.

18.1 \maketitle

The \maketitle command generates a title on a separate title page—except in the article class, where the title is placed at the top of the first page. Information used to produce the title is obtained from the following declarations:

\author{name \and name2}

The \author command declares the document author(s), where the argument is a list of authors separated by \and commands. Use \\ to separate lines within a single author’s entry—for example, to give the author’s institution or address.

\date{text}

The \date command declares text to be the document’s date. With no \date command, the current date (see Section 21.6 [\today], page 69) is used.

\thanks{text}

The \thanks command produces a \footnote to the title, usually used for credit acknowledgements.

\title{text}

The \title command declares text to be the title of the document. Use \\ to force a line break, as usual.

18.2 \pagenumbering

Synopsis:

     \pagenumbering{style}

Specifies the style of page numbers, according to style:

arabic arabic numerals

roman lowercase Roman numerals

Roman uppercase Roman numerals

alph lowercase letters

Alph uppercase letters

18.3 \pagestyle

Synopsis:

     \pagestyle{style}

The \pagestyle command specifies how the headers and footers are typeset from the current page onwards. Values for style:

plain Just a plain page number. empty Empty headers and footers, e.g., no page numbers.

headings Put running headers on each page. The document style specifies what goes in the headers.

myheadings

Custom headers, specified via the \markboth or the \markright commands.

Here are the descriptions of \markboth and \markright:

\markboth{left}{right}

Sets both the left and the right heading. A “left-hand heading” (left) is generated by the last \markboth command before the end of the page, while a “right-hand heading” (right is generated by the first \markboth or \markright that comes on the page if there is one, otherwise by the last one before the page.

\markright{right}

Sets the right heading, leaving the left heading unchanged.

18.4 \thispagestyle{style}

The \thispagestyle command works in the same manner as the \pagestyle command (see previous section) except that it changes to style for the current page only.

19 Spaces

L^{A}TEX has many ways to produce white (or filled) space.

Another space-producing command is \, to produce a “thin” space (usually 1/6 quad). It can be used in text mode, but is more often useful in math mode (see Section 16.5 [Spacing in math mode], page 56).

19.1 \hspace

Synopsis:

     \hspace[*]{length}

The \hspace command adds horizontal space. The length argument can be expressed in any terms that L^{A}TEX understands: points, inches, etc. It is a rubber length. You can add both negative and positive space with an \hspace command; adding negative space is like backspacing.

L^{A}TEX normally removes horizontal space that comes at the beginning or end of a line. To preserve this space, use the optional * form.

19.2 \hfill

The \hfill fill command produces a “rubber length” which has no natural space but can stretch or shrink horizontally as far as needed.

The \fill parameter is the rubber length itself (technically, the glue value ‘0pt

plus1fill’); thus, \hspace\fill is equivalent to \hfill.
19.3 \SPACE

The \ (space) command produces a normal interword space. It’s useful after punctuation which shouldn’t end a sentence. For example Knuth's article in Proc.\ Amer.\ Math\. Soc.\ is fundamental. It is also often used after control sequences, as in \TeX\ is a nice

system.

In normal circumstances, \TAB and \NEWLINE are equivalent to \ .

19.4 \@

The \@ command makes the following punctuation character end a sentence even if it normally would not. This is typically used after a capital letter. Here are side-by-side examples with and without \@:

     ... in C\@.  Pascal, though ...
     ... in C.  Pascal, though ...

produces … in C. Pascal, though . . . . . . in C. Pascal, though …

19.5 \thinspace

\thinspace produces an unbreakable and unstretchable space that is 1/6 of an em. This is the proper space to use in nested quotes, as in ’ ”.

19.6 \/

The \/ command produces an italic correction. This is a small space defined by the font designer for a given character, to avoid the character colliding with whatever follows. The italic f character typically has a large italic correction value.

If the following character is a period or comma, it’s not necessary to insert an italic correction, since those punctuation symbols have a very small height. However, with semicolons or colons, as well as normal letters, it can help. Compare f : f ; with f: f;.

Despite the name, roman characters can also have an italic correction. Compare pdfTEX with pdfTEX.

19.7 \hrulefill

The \hrulefill fill command produces a “rubber length” which can stretch or shrink horizontally. It will be filled with a horizontal rule.

19.8 \dotfill

The \dotfill command produces a “rubber length” that fills with dots instead of just white space.

19.9 \addvspace
\addvspace{length}

The \addvspace command normally adds a vertical space of height length. However, if vertical space has already been added to the same point in the output by a previous \addvspace command, then this command will not add more space than needed to make the natural length of the total vertical space equal to length.

19.10 \bigskip \medskip \smallskip

These commands produce a given amount of space.

\bigskip The same as \vspace{bigskipamount}, ordinarily about one line space (with stretch and shrink).

\medskip The same as \vspace{medskipamount}, ordinarily about half of a line space (with stretch and shrink).

\smallskip

The same as \vspace{smallskipamount}, ordinarily about a quarter of a line space (with stretch and shrink).

The \...amount parameters are determined by the document class.

19.11 \vfill

The \vfill fill command produces a rubber length (glue) which can stretch or shrink vertically as far as needed. It’s equivalent to \vspace{\fill} (see Section 19.2 [\hfill], page 61).

19.12 \vspace[*]{length}

Synopsis:

     \vspace[*]{length}

The \vspace command adds the vertical space length, i.e., a rubber length. length can be negative or positive.

Ordinarily, L^{A}TEX removes vertical space added by \vspace at the top or bottom of a page. With the optional * argument, the space is not removed.

20 Boxes

All the predefined length parameters (see Section 12.6 [Predefined lengths], page 40) can be used in the arguments of the box-making commands.

20.1 \mbox{text}

The \mbox command creates a box just wide enough to hold the text created by its argument. The text is not broken into lines, so it can be used to prevent hyphenation.

20.2 \fbox and \framebox

Synopses:

     \fbox{text}
     \framebox[width][position]{text}

The \fbox and \framebox commands are like \mbox, except that they put a frame around the outside of the box being created.

In addition, the \framebox command allows for explicit specification of the box width with the optional width argument (a dimension), and positioning with the optional position argument.

Both commands produce a rule of thickness \fboxrule (default ‘.4pt’), and leave a space of \fboxsep (default ‘3pt’) between the rule and the contents of the box.

See Section 8.18.3 [\framebox (picture)], page 23, for the \framebox command in the picture environment.

20.3 lrbox
\begin{lrbox}{cmd} text \end{lrbox}

This is the environment form of \sbox.

The text inside the environment is saved in the box cmd, which must have been declared

with \newsavebox.
20.4 \makebox

Synopsis:

     \makebox[width][position]{text}

The \makebox command creates a box just wide enough to contain the text specified. The width of the box is specified by the optional width argument. The position of the text within the box is determined by the optional position argument, which may take the following values:

c Centered (default).

l Flush left.

r Flush right.

s Stretch (justify) across entire width; text must contain stretchable space for this to work.

\makebox is also used within the picture environment see Section 8.18.2 [\makebox (picture)], page 23.

20.5 \parbox

Synopsis:

     \parbox[position][height][inner-pos]{width}{text}

The \parbox command produces a box whose contents are created in paragraph mode. It should be used to make a box small pieces of text, with nothing fancy inside. In particular, you shouldn’t use any paragraph-making environments inside a \parbox argument. For larger pieces of text, including ones containing a paragraph-making environment, you should use a minipage environment (see Section 8.17 [minipage], page 21).

\parbox has two mandatory arguments:

width the width of the parbox;

text the text that goes inside the parbox.

The optional position argument allows you to align either the top or bottom line in the parbox with the baseline of the surrounding text (default is top).

The optional height argument overrides the natural height of the box.

The inner-pos argument controls the placement of the text inside the box, as follows; if it is not specified, position is used.

t text is placed at the top of the box.

c text is centered in the box.

b text is placed at the bottom of the box.

s stretch vertically; the text must contain vertically stretchable space for this to work.

20.6 \raisebox

Synopsis:

     \raisebox{distance}[height][depth]{text}

The \raisebox command raises or lowers text. The first mandatory argument specifies how high text is to be raised (or lowered if it is a negative amount). text itself is processed in LR mode.

The optional arguments height and depth are dimensions. If they are specified, L^{A}TEX treats text as extending a certain distance above the baseline (height) or below (depth), ignoring its natural height and depth.

20.7 \savebox

Synopsis:

     \savebox{\boxcmd}[width][pos]{text}

This command typeset text in a box just as with \makebox (see Section 20.4 [\makebox], page 64), except that instead of printing the resulting box, it saves it in the box labeled \boxcmd, which must have been declared with \newsavebox (see Section 10.4 [\newsavebox], page 36).

20.8 \sbox{\boxcmd}{text}

Synopsis:

     \sbox{\boxcmd}{text}

\sbox types text in a box just as with \mbox (see Section 20.1 [\mbox], page 64) except that instead of the resulting box being included in the normal output, it is saved in the box labeled \boxcmd. \boxcmd must have been previously declared with \newsavebox (see Section 10.4 [\newsavebox], page 36).

20.9 \usebox{\boxcmd

Synopsis:

     \usebox{\boxcmd}

\usebox producesthe box most recently saved in the bin \boxcmd by a \savebox command (see Section 20.7 [\savebox], page 65).

21 Special insertions

L^{A}TEX provides commands for inserting characters that have a special meaning do not correspond to simple characters you can type.

21.1 Reserved characters

The following characters play a special role in L^{A}TEX and are called “reserved characters” or “special characters”.

     # $ % & ~ _^\ { }

Whenever you write one of these characters into your file, L^{A}TEX will do something special. If you simply want the character to be printed as itself, include a \ in front of the character. For example, \$ will produce $ in your output.

One exception to this rule is \ itself, because \\ has its own special (context-dependent) meaning. A roman \ is produced by typing $\backslash$ in your file, and a typewriter \ is produced by using ‘\’ in a verbatim command (see Section 8.27 [verbatim], page 32).

Also, \~ and \^place tilde and circumflex accents over the following letter, as in õ and ô (see Section 21.3 [Accents], page 68); to get a standalone ~ or ^, you can again use a verbatim command.

Finally, you can access any character of the current font once you know its number by using the \symbol command. For example, the visible space character used in the \verb* command has the code decimal 32, so it can be typed as \symbol{32}.

You can also specify octal numbers with ' or hexadecimal numbers with ", so the previous example could also be written as \symbol{'40} or \symbol{"20}.

21.2 Text symbols

L^{A}TEX provides commands to generate a number of non-letter symbols in running text.

\copyright

The copyright symbol, \begin{matrix}c \\ ◯ \end{matrix}.

\dag The dagger symbol (in text).

\ddag The double dagger symbol (in text).

\LaTeX The L^{A}TEX logo.

\ldots An ellipsis (three dots at the baseline): ‘…’. This command also works in math mode.

\lq Left (opening) quote: ‘.

\P Paragraph sign (pilcrow).

\pounds English pounds sterling.

\rq Right (closing) quote: ’.

\S Section symbol.

\TeX The TEX logo.

21.3 Accents

L^{A}TEX has wide support for many of the world’s scripts and languages, through the babel package and related support. This section does not attempt to cover all that support. It merely the core L^{A}TEX commands for creating accented characters. \" Produces an umlaut, as in ö. \' Produces an acute accent, as in ó. In the tabbing environment, pushes current

column to the right of the previous column (see Section 8.21 [tabbing], page 26).

\. Produces a dot accent over the following, as in ȯ. \= Produces a macron (overbar) accent over the following, as in ō. \^Produces a circumflex (hat) accent over the following, as in ô. \` Produces a grave accent over the following, as in ò. In the tabbing environment,

move following text to the right margin (see Section 8.21 [tabbing], page 26).

\~ Produces a tilde accent over the following, as in ñ. \b Produces a bar accent under the following, as in ō. \bar Produces a macron accent over the following, as in ō. \c Produces a cedilla accent under the following, as in \d Produces a dot accent under the following, as in o.. \H Produces a long Hungarian umlaut accent over the following, as in ő. \i Produces a dotless i, as in ‘ı’. \j Produces a dotless j, as in ‘ȷ’. \t Produces a tie-after accent, as in ‘oo’. \u Produces a breve accent, as in ‘ŏ’. \v Produces a háček (check) accent, as in ‘ǒ’.

21.4 Non-English characters

Here are the basic L^{A}TEX commands for inserting characters commonly used in languages other than English.

\aa

\AA å and Å.

\ae

\AE æ and Æ.

\l

\L ł and Ł.

\o

\O ø and Ø.

\oe

\OE œ and Œ.

\ss  ß.

21.5 \rule

Synopsis:

     \rule[raise]{width}{thickness}

The \rule command produces rules, that is, lines or rectangles. The arguments are: raise How high to raise the rule (optional). width The length of the rule (mandatory). thickness The thickness of the rule (mandatory).

21.6 \today

The \today command produces today’s date, in the format ‘month dd, yyyy’; for example, ‘July 4, 1976’. It uses the predefined counters \day, \month, and \year (see Section 11.8 [\day \month \year], page 39) to do this. It is not updated as the program runs.

The datetime package, among others, can produce a wide variety of other date formats.

22 Splitting the input

A large document requires a lot of input. Rather than putting the whole input in a single large file, it’s more efficient to split it into several smaller ones. Regardless of how many separate files you use, there is one that is the root file; it is the one whose name you type when you run L^{A}TEX.

22.1 \include

Synopsis:

     \include{file}

If no \includeonly command is present, the \include command executes \clearpage to start a new page (see Section 14.2 [\clearpage], page 43), then reads file, then does

another \clearpage.

Given an \includeonly command, the \include actions are only run if file is listed as an argument to \includeonly. See the next section.

The \include command may not appear in the preamble or in a file read by another

\include command.
22.2 \includeonly

Synopsis:

     \includeonly{file1,file2,...}

The \includeonly command controls which files will be read by subsequent \include commands. The list of filenames is comma-separated. Each file must exactly match a filename specified in a \include command for the selection to be effective.

This command can only appear in the preamble.

22.3 \input

Synopsis:

     \input{file}

The \input command causes the specified file to be read and processed, as if its contents had been inserted in the current file at that point.

If file does not end in ‘.tex’ (e.g., ‘foo’ or ‘foo.bar’), it is first tried with that extension (‘foo.tex’ or ‘foo.bar.tex’). If that is not found, the original file is tried (‘foo’ or

‘foo.bar’).

23 Front/back matter

23.1 Tables of contents

A table of contents is produced with the \tableofcontents command. You put the command right where you want the table of contents to go; L^{A}TEX does the rest for you. A previous run must have generated a ‘.toc’ file.

The \tableofcontents command produces a heading, but it does not automatically start a new page. If you want a new page after the table of contents, write a \newpage command after the \tableofcontents command.

The analagous commands \listoffigures and \listoftables produce a list of figures and a list of tables, respectively. Everything works exactly the same as for the table of contents.

The command \nofiles overrides these commands, and prevents any of these lists from being generated.

23.1.1 \addcontentsline

The \addcontentsline command adds an entry to the specified list or table where:

ext The extension of the file on which information is to be written, typically one of: toc (table of contents), lof (list of figures), or lot (list of tables).

unit The name of the sectional unit being added, typically one of the following, matching the value of the ext argument:

toc The name of the sectional unit: part, chapter, section,

                    subsection, subsubsection.

lof For the list of figures.

lot For the list of tables.

entry The actual text of the entry.

What is written to the ‘.ext’ file is the command \contentsline{unit}{name}.

23.1.2 \addtocontents

The \addtocontents command adds text (or formatting commands) directly to the ‘.’ext file that generates the table of contents or lists of figures or tables.

ext The extension of the file on which information is to be written: ‘toc’ (table of contents), ‘lof’ (list of figures), or ‘lot’ (list of tables).

text The text to be written.

23.2 Glossaries

The command \makeglossary enables creating glossaries.

The command \glossary{text} writes a glossary entry for text to an auxiliary file with the ‘.glo’ extension.

Specifically, what gets written is the command \glossaryentry{text}{pageno}, where pageno is the current \thepage value.

The glossary package on CTAN provides support for fancier glossaries.

23.3 Indexes

The command \makeindex enables creating indexes.

The command \index{text} writes a glossary entry for text to an auxiliary file with the ‘.idx’ extension.

Specifically, what gets written is the command \indexentry{text}{pageno}, where pageno is the current \thepage value.

The rubber length \indexspace is inserted before each new letter in the index; its default

value is ‘10pt plus5pt minus3pt’.

The ‘.idx’ file can then be sorted with an external command, typically makeindex or xindy, resulting in a ‘.ind’ file, which can then be used to typeset the index.

24 Letters

You can use L^{A}TEX to typeset letters, both personal and business. The letter document class is designed to make a number of letters at once, although you can make just one if you so desire.

Your ‘.tex’ source file has the same minimum commands as the other document classes, i.e., you must have the following commands as a minimum:

      \documentclass{letter}
      \begin{document}
       ... letters ...
      \end{document}

Each letter is a letter environment, whose argument is the name and address of the recipient. For example, you might have:

      \begin{letter}{Mr. Joe Smith\\ 2345 Princess St.
           \\ Edinburgh, EH1 1AA}
        ...
      \end{letter}

The letter itself begins with the \opening command. The text of the letter follows. It is typed as ordinary L^{A}TEX input. Commands that make no sense in a letter, like \chapter, do not work. The letter closes with a \closing command.

After the closing, you can have additional material. The \cc command produces the usual “cc: …”. There’s also a similar \encl command for a list of enclosures. With both these commands, use \\ to separate the items.

These commands are used with the letter class.

24.1 \address{return-address}

The \address specifies the return address of a letter, as it should appear on the letter and the envelope. Separate lines of the address should be separated by \\ commands.

If you do not make an \address declaration, then the letter will be formatted for copying onto your organisation’s standard letterhead. (See Chapter 1 [Overview], page 2, for details on your local implementation). If you give an \address declaration, then the letter will be formatted as a personal letter.

24.2 \cc

Synopsis:

     \cc{name1\\name2}

Produce a list of names the letter was copied to. Each name is printed on a separate line.

24.3 \closing

Synopsis:

     \closing{text}

A letter closes with a \closing command, for example,

     \closing{Best Regards,}

24.4 \encl

Synopsis:

     \encl{line1\\line2}

Declare a list of one more enclosures.

24.5 \location
\location{address}

This modifies your organisation’s standard address. This only appears if the firstpage pagestyle is selected.

24.6 \makelabels
\makelabels{number}

If you issue this command in the preamble, L^{A}TEX will create a sheet of address labels. This sheet will be output before the letters.

24.7 \name
\name{June Davenport}

Your name, used for printing on the envelope together with the return address.

24.8 \opening{text}

Synopsis:

     \opening{text}

A letter begins with the \opening command. The mandatory argument, text, is whatever text you wish to start your letter. For instance:

     \opening{Dear Joe,}

24.9 \ps

Use the \ps command to start a postscript in a letter, after \closing.

24.10 \signature{text}

Your name, as it should appear at the end of the letter underneath the space for your signature. \\ starts a new line within text as usual.

24.11 \startbreaks
\startbreaks

Used after a \stopbreaks command to allow page breaks again.

24.12 \stopbreaks
\stopbreaks

Inhibit page breaks until a \startbreaks command occurs.

Chapter 24: Letters 75

24.13 \telephone
\telephone{number}

This is your telephone number. This only appears if the firstpage pagestyle is selected.

25 Terminal Input/Output

25.1 \typein[cmd]{msg}

Synopsis:

     \typein[\cmd]{msg}

\typein prints msg on the terminal and causes L^{A}TEX to stop and wait for you to type a line of input, ending with return. If the optional \cmd argument is omitted, the typed input is processed as if it had been included in the input file in place of the \typein command. If the \cmd argument is present, it must be a command name. This command name is then defined or redefined to be the typed input.

25.2 \typeout{msg}

Synopsis:

     \typeout{msg}

Prints msg on the terminal and in the log file. Commands in msg that are defined with \newcommand or \renewcommand (among others) are replaced by their definitions before being printed.

L^{A}TEX’s usual rules for treating multiple spaces as a single space and ignoring spaces after a command name apply to msg. A \space command in msg causes a single space to be printed, independent of surrounding spaces. A ^^J in msg prints a newline.

26 Command Line

The input file specification indicates the file to be formatted; TEX uses ‘.tex’ as a default file extension. If you omit the input file entirely, TEX accepts input from the terminal. You specify command options by supplying a string as a parameter to the command; e.g.

     latex '\nonstopmode\input foo.tex'

will process ‘foo.tex’ without pausing after every error.

If L^{A}TEX stops in the middle of the document and gives you a ‘*’ prompt, it is waiting for input. You can type \stop (and return) and it will prematurely end the document.

Appendix A Document templates

Although not reference material, perhaps these document templates will be useful.

A.1 book template

\documentclass{book}
\title{Book Class Template}
\author{Alex Author}
\begin{document}
\maketitle
\chapter{First}
Some text.
\chapter{Second}
Some other text.
\section{A subtopic}
The end.
\end{document}

A.2 beamer template

The beamer class creates slides presentations.

\documentclass{beamer}
\title{Beamer Class template}
\author{Alex Author}
\date{July 31, 2007}
\begin{document}
\maketitle
% without [fragile], any {verbatim} code gets mysterious errors.
\begin{frame}[fragile]
 \frametitle{First Slide}
\begin{verbatim}
  This is \verbatim!
\end{verbatim}
\end{frame}
\end{document}

A.3 tugboat template

TUGboat is the journal of the TEX Users Group, http://tug.org/TUGboat.

\documentclass{ltugboat}  % ltugproc for proceedings
\usepackage{graphicx}
\usepackage{ifpdf}
\ifpdf
  \usepackage[breaklinks,colorlinks,linkcolor=black,citecolor=black,
              urlcolor=black]{hyperref}
\else
  \usepackage{url}
\fi
\begin{document}
\title{Example \TUB\ Regular Article}
% repeat info for each author.
\author{First Last}
\address{Street Address \\ Town, Postal \\ Country}
\netaddress{user (at) example dot org}
\personalURL{http://example.org/~user/}
\maketitle
% The abstract comes after \maketitle in ltugboat.
\begin{abstract}
This is an example article for a regular \TUB{} issue.
\end{abstract}
\section{Introduction}
This is an example article for \TUB, from
\url{http://tug.org/TUGboat/location.html}.
We recommend the graphicx package for image inclusions, and the hyperref
package for active url's (in the \acro{PDF} output).  \TUB\ is produced
using \acro{PDF} files exclusively, nowadays.
The \texttt{ltug*} classes provide these abbreviations, among many others:
{\small
\begin{verbatim}
\AllTeX \AMS \AmS \AmSLaTeX \AmSTeX \aw \AW
\BibTeX \CandT \CTAN \DTD \DVItoVDU \HTML
\ISBN \ISSN \JTeX \JoT \LAMSTeX \LaTeXe

\Mc \mf \MFB \mtex \pcMF \PCTeX \pcTeX \Pas
\PiC \PiCTeX \plain \POBox \PS
\SC \SGML \SliTeX \TANGLE \TB \TP \TUB \TUG
\tug \UG \UNIX \VAX \VorTeX \XeT \WEB \WEAVE
\Dash \dash \vellipsis \bull \cents \Dag
\careof \thinskip
\acro{FRED} -> {\sc fred}  % please use!
\cs{fred}   -> \fred
\env{fred}  -> \begin{fred}
\meta{fred} -> <fred>
\nth{n}     -> 1st, 2nd, ...
\sfrac{3/4} -> 3/4
\booktitle{Book of Fred}
\end{verbatim}
}
For more information, see the ltubguid document at:
\url{http://mirror.ctan.org/macros/latex/contrib/tugboat}
Email \verb|tugboat@tug.org| if problems or questions.

\bibliographystyle{plain}  % we recommend the plain bibliography style
\nocite{book-minimal}      % just making the bibliography non-empty
\bibliography{xampl}       % xampl.bib comes with BibTeX

\makesignature  % not in ltugproc
\end{document}

Concept Index 81

Concept Index

*

‘*’ prompt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 *-form of sectioning commands . . . . . . . . . . . . . . . . 12

.

‘.glo’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 ‘.idx’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 ‘.ind’ file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

A

abstracts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 accents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 accents, mathematical . . . . . . . . . . . . . . . . . . . . . . . . . 55 accessing any character of a font . . . . . . . . . . . . . . . 67 acute accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 acute accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 ae ligature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 aligning Equations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 alignment via tabbing . . . . . . . . . . . . . . . . . . . . . . . . . 26 appendix, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 aring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 arrays, math. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 author, for titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

B

bar-over accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 bar-under accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 bibliography, creating (automatically) . . . . . . . . . . 31 bibliography, creating (manually) . . . . . . . . . . . . . . 30 bibTEX, using . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 black boxes, omitting . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 bold font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 bold typewriter, avoiding . . . . . . . . . . . . . . . . . . . . . . 15 boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 breaking lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 breaking pages . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 breve accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 breve accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 bullet symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 bulleted lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19

C

calligraphic letters for math . . . . . . . . . . . . . . . . . . . . . 7 cc list, in letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 cedilla accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 centering text, declaration for . . . . . . . . . . . . . . . . . . 15 centering text, environment for . . . . . . . . . . . . . . . . 15 characters, accented . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 characters, non-English . . . . . . . . . . . . . . . . . . . . . . . . 68 characters, reserved. . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 check accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 check accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 circumflex accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 circumflex accent, math. . . . . . . . . . . . . . . . . . . . . . . . 55 class options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 classes of documents . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 closing letters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 closing quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 code, typesetting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 command Line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77 commands, defining new ones . . . . . . . . . . . . . . . . . . 35 computer programs, typesetting. . . . . . . . . . . . . . . . 32 copyright symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 counters, a list of . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 counters, defining new . . . . . . . . . . . . . . . . . . . . . . . . . 35 counters, getting value of . . . . . . . . . . . . . . . . . . . . . . 38 counters, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 creating letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 creating pictures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 creating tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 credit footnote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 cross references . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 cross referencing with page number . . . . . . . . . . . . 13 cross referencing, symbolic . . . . . . . . . . . . . . . . . . . . . 13

D

dagger, in text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 date, for titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 datetime package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 defining a new command. . . . . . . . . . . . . . . . . . . . . . . 35 defining new environments . . . . . . . . . . . . . . . . . . . . . 36 defining new fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 defining new theorems . . . . . . . . . . . . . . . . . . . . . . . . . 36 definitions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35 description lists, creating . . . . . . . . . . . . . . . . . . . . . . 15 discretionary multiplication . . . . . . . . . . . . . . . . . . . . 56 displaying quoted text with paragraph indentation

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 displaying quoted text without paragraph

indentation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 document class options . . . . . . . . . . . . . . . . . . . . . . . . . 4 document classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 document templates . . . . . . . . . . . . . . . . . . . . . . . . . . . 78 dot accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 dot over accent, math. . . . . . . . . . . . . . . . . . . . . . . . . . 55 dot-over accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 dot-under accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 dotless i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 dotless i, math. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 dotless j . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 dotless j, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 double dagger, in text . . . . . . . . . . . . . . . . . . . . . . . . . 67 double dot accent, math . . . . . . . . . . . . . . . . . . . . . . . 55

E

ellipsis. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 emphasis . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7 enclosure list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 ending & starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 enlarge current page . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 environments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 environments, defining . . . . . . . . . . . . . . . . . . . . . . . . . 36 equation number, cross referencing . . . . . . . . . . . . . 13 equation numbers, omitting . . . . . . . . . . . . . . . . . . . . 17 equations, aligning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 equations, environment for . . . . . . . . . . . . . . . . . . . . . 17 es-zet German letter . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 exponent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

F

figure number, cross referencing . . . . . . . . . . . . . . . . 13 figures, footnotes in. . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 figures, inserting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 fixed-width font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 flushing floats and starting a page. . . . . . . . . . . . . . 43 font commands, low-level . . . . . . . . . . . . . . . . . . . . . . . 8 font sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 font styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 fonts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 fonts, new commands for . . . . . . . . . . . . . . . . . . . . . . 37 footer style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 footer, parameters for. . . . . . . . . . . . . . . . . . . . . . . . . . 11 footnote number, cross referencing . . . . . . . . . . . . . 13 footnote parameters . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 footnotes in figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 footnotes, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 formulas, environment for. . . . . . . . . . . . . . . . . . . . . . 17 formulas, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 fragile commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 functions, math. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

G

global options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 5 glossaries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 grave accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 grave accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 greek letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

H

hacek accent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 háček accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 hat accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 hat accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 header style . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 header, parameters for . . . . . . . . . . . . . . . . . . . . . . . . . 11 hungarian umlaut accent. . . . . . . . . . . . . . . . . . . . . . . 68 hyphenation, defining . . . . . . . . . . . . . . . . . . . . . . . . . . 42 hyphenation, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 hyphenation, preventing . . . . . . . . . . . . . . . . . . . . . . . 64

I

in-line formulas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 indent, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 indent, suppressing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 indentation of paragraphs, in minipage . . . . . . . . . 21 indexes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 input file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 input/Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 inserting figures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

J

justification, ragged left. . . . . . . . . . . . . . . . . . . . . . . . 19 justification, ragged right . . . . . . . . . . . . . . . . . . . . . . 18

K

Knuth, Donald E. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

L

labelled lists, creating. . . . . . . . . . . . . . . . . . . . . . . . . . 15 Lamport, Leslie. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 L^{A}TEX logo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 L^{A}TEX overview . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 L^{A}TEX Project team. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 layout commands . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 layout, page parameters for . . . . . . . . . . . . . . . . . . . . 11 left quote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 left-justifying text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 left-justifying text, environment for . . . . . . . . . . . . 18 left-to-right mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 lengths, adding to . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 lengths, defining and using. . . . . . . . . . . . . . . . . . . . . 40 lengths, defining new . . . . . . . . . . . . . . . . . . . . . . . . . . 35 lengths, predefined. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 lengths, setting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 letters, accented . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 letters, ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 letters, non-English . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 letters, starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 line break, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 line breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 line breaks, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 line breaks, preventing . . . . . . . . . . . . . . . . . . . . . . . . . 42 lines in tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 lining text up in columns using tab stops. . . . . . . 26 lining text up in tables . . . . . . . . . . . . . . . . . . . . . . . . 28 list items, specifying counter . . . . . . . . . . . . . . . . . . . 38 lists of items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 lists of items, generic . . . . . . . . . . . . . . . . . . . . . . . . . . 21 lists of items, numbered. . . . . . . . . . . . . . . . . . . . . . . . 16 loading additional packages . . . . . . . . . . . . . . . . . . . . . 5 logo, L^{A}TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 logo, TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 low-level font commands . . . . . . . . . . . . . . . . . . . . . . . . 8 lR mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58

M

macron accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 makeindex program. . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 making a title page . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 making paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 marginal notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 math accents . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 math formulas . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 math functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 math Miscellany . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 math mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 math mode, entering. . . . . . . . . . . . . . . . . . . . . . . . . . . 46 math mode, spacing . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 math symbols . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 minipage, creating a . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 monospace font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 moving arguments . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 multicolumn text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 multiplication symbol, discretionary line break

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56

N

nested \include, not allowed . . . . . . . . . . . . . . . . . . 70 new commands, defining . . . . . . . . . . . . . . . . . . . . . . . 35 new line, output as input . . . . . . . . . . . . . . . . . . . . . . 41 new line, starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 new line, starting (paragraph mode) . . . . . . . . . . . 41 new page, starting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 non-English characters. . . . . . . . . . . . . . . . . . . . . . . . . 68 notes in the margin . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 null delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 numbered items, specifying counter . . . . . . . . . . . . 38

O

oblique font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 oe ligature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 one-column output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 opening quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 options, document class . . . . . . . . . . . . . . . . . . . . . . . . . 4 options, global . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 oslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 overbar accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 overdot accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . 55 overview of L^{A}TEX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

P

packages, loading . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 page break, forcing . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 page break, preventing. . . . . . . . . . . . . . . . . . . . . . . . . 43 page breaking . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 page layout parameters . . . . . . . . . . . . . . . . . . . . . . . . 11 page number, cross referencing . . . . . . . . . . . . . . . . . 13 page numbering style . . . . . . . . . . . . . . . . . . . . . . . . . . 59 page styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 paragraph indentation, in minipage . . . . . . . . . . . . 21 paragraph indentations in quoted text . . . . . . . . . 25 paragraph indentations in quoted text, omitting

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 paragraph mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 58 paragraph sign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 parameters, page layout . . . . . . . . . . . . . . . . . . . . . . . 11 pictures, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 pilcrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 poetry, an environment for. . . . . . . . . . . . . . . . . . . . . 32 polish l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 postscript, in letters . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 pounds symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 preamble, defined . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 predefined lengths . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 prompt, ‘*’ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77

Q

quoted text with paragraph indentation, displaying

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 quoted text without paragraph indentation,

displaying. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

R

ragged left text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 ragged left text, environment for . . . . . . . . . . . . . . . 19 ragged right text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 ragged right text, environment for . . . . . . . . . . . . . 18 redefining environments. . . . . . . . . . . . . . . . . . . . . . . . 36 remarks in the margin . . . . . . . . . . . . . . . . . . . . . . . . . 44 reserved characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 right quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 right-justifying text . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 right-justifying text, environment for . . . . . . . . . . . 19 robust commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 roman font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 running header and footer . . . . . . . . . . . . . . . . . . . . . 11 running header and footer style . . . . . . . . . . . . . . . . 59

S

sans serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 script letters for math . . . . . . . . . . . . . . . . . . . . . . . . . . 7 section number, cross referencing . . . . . . . . . . . . . . 13 section symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 sectioning . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 setspace package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 setting counters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 sharp S letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 simulating typed text . . . . . . . . . . . . . . . . . . . . . . . . . . 32 sizes of text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 slanted font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 small caps font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 space, inserting vertical . . . . . . . . . . . . . . . . . . . . . . . . 62 spaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 spacing within math mode . . . . . . . . . . . . . . . . . . . . . 56 special characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 splitting the input file . . . . . . . . . . . . . . . . . . . . . . . . . 70 starting & ending . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 starting a new page. . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 starting a new page and clearing floats . . . . . . . . . 43 starting on a right-hand page . . . . . . . . . . . . . . . . . . 43 sterling symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 stretch, omitting vertical. . . . . . . . . . . . . . . . . . . . . . . 11 styles of text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 styles, page . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 subscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 superscript . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 symbols, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

T

tab stops, using. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 table of contents entry, manually adding . . . . . . . 71 table of contents, creating . . . . . . . . . . . . . . . . . . . . . 71 tables, creating . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 terminal Input/Output . . . . . . . . . . . . . . . . . . . . . . . . 76 TEX logo . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 text symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 thanks, for titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 theorems, defining . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 theorems, typesetting . . . . . . . . . . . . . . . . . . . . . . . . . . 31 tie-after accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 tilde accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 tilde accent, math . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 title pages, creating. . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 title, for titlepage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 titles, making. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 two-column output . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 typed text, simulating . . . . . . . . . . . . . . . . . . . . . . . . . 32 typeface sizes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 typeface styles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 typefaces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 typewriter font . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 typewriter labels in lists . . . . . . . . . . . . . . . . . . . . . . . 15

U

umlaut accent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 unordered lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 using BibTEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31

V

variables, a list of. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 vector symbol, math . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 verbatim text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 verbatim text, inline . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 vertical space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 vertical space before paragraphs . . . . . . . . . . . . . . . 44 visible space . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32

W

wide hat accent, math . . . . . . . . . . . . . . . . . . . . . . . . . 55 wide tile accent, math . . . . . . . . . . . . . . . . . . . . . . . . . 55

X

xindy program . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

Command Index 85

Command Index

$

$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

@

@{...} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

^

^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

\

\" (umlaut accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \# . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \$ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \% . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \& . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \' (acute accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \' (tabbing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \( . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \, . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \- (hyphenation) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 \. (dot-over accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \/ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \: . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \; . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \< . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \= (macron accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \= (tabbing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \> . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26, 56 \> (tabbing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \@ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 \[ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \^. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \^(circumflex accent) . . . . . . . . . . . . . . . . . . . . . . . . . 68 \_. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \` (grave accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \` (tabbing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \{ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \\ (for \shortstack objects) . . . . . . . . . . . . . . . . . . 25 \\ (for array) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 \\ (for center). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 \\ (for eqnarray) . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 \\ (for flushright) . . . . . . . . . . . . . . . . . . . . . . . . . . 19 \\ for \author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \\ for \title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \\ for flushleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \\ for letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 \\ for tabular. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 \\ for verse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 \\ force line break . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 \\ tabbing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 26 \\* (for eqnarray) . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 \| . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \~ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \~ (tilde accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \a (tabbing) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 \a' (acute accent in tabbing) . . . . . . . . . . . . . . . . . . 27 \a= (macron accent in tabbing) . . . . . . . . . . . . . . . . 27 \a` (grave accent in tabbing) . . . . . . . . . . . . . . . . . . 27 \aa (å) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \AA (Å) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \acute . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55

\addcontentsline{ext}{unit}{text} . . . . . . . . 71

\address . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73

\addtocontents{ext}{text} . . . . . . . . . . . . . . . . . . 71

\addtocounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 \addtolength . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 \addvspace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \ae (æ) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \AE (Æ) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \aleph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \alph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \Alph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \Alph example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 \alpha . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \amalg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \and for \author . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \angle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \appendix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 \approx . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \arabic . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \arccos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \arcsin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \arctan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \arg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \arraycolsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 \arrayrulewidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 \arraystretch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 \ast . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \asymp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47

\author{name \and name2} . . . . . . . . . . . . . . . . . . . . 59

\b (bar-under accent). . . . . . . . . . . . . . . . . . . . . . . . . . 68 \backslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47, 67 \bar (macron, bar-over accent) . . . . . . . . . . . . . . . . 68 \baselineskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \baselinestretch . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \begin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 \beta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \bfseries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \bibitem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \bibliography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 \bibliographystyle . . . . . . . . . . . . . . . . . . . . . . . . . . 31 \bigcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigcirc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigcup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigodot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigoplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigotimes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \bigskipamount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \bigsqcup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigtriangledown. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigtriangleup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \biguplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bigwedge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \boldmath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \bot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \bottomfraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \bowtie . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \Box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \breve . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \bullet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \c (cedilla accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \cal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \cap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \caption . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \cc. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 \cdot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \cdots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \centering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 \chapter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 \check . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \chi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \circ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \circle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 \cite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31 \cleardoublepage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 \clearpage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43 \cline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \closing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 73 \clubsuit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \columnsep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 \columnseprule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 \columnwidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 \cong . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \contentsline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \coprod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 \copyright. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \cos . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \cosh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \cot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \coth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \csc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

\cup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\d (dot-under accent). . . . . . . . . . . . . . . . . . . . . . . . . . 68
\dag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\dagger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\dashbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23
\dashv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\date{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
\day . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
\dblfloatpagefraction . . . . . . . . . . . . . . . . . . . . . . . 10
\dblfloatsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
\dbltextfloatsep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
\dbltopfraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
\ddag . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\ddagger . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\ddot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\ddots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
\deg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\Delta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\depth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
\det . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\diamond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\Diamond . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\diamondsuit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\displaystyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46
\div . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\documentclass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
\dot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\doteq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\dotfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
\doublerulesep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29
\downarrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\Downarrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\ell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\em . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\emph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\emptyset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\encl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\end . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14
\enlargethispage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43
\enumi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\enumii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\enumiii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\enumiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\epsilon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\equiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\eta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\evensidemargin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
\exists . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\exp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\fbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64
\fboxrule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 64
\fboxsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 64
\fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
\flat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48
\floatpagefraction . . . . . . . . . . . . . . . . . . . . . . . . . . 18
\floatsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

\flushbottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \fnsymbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \fontencoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \fontfamily. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \fontseries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \fontshape. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \fontsize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 \footnote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 \footnotemark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 \footnoterule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 \footnotesep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 \footnotesize . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \footnotetext . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34 \footskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \forall . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \frac . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \frac{num}{den} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \frame . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 \framebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23, 64 \frown . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \fussy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 \gamma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \Gamma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \gcd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \ge. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \geq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \gets . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \gg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \glossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \glossaryentry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \grave . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \H (Hungarian umlaut accent) . . . . . . . . . . . . . . . . . 68 \hat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \hbar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \headheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \headsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \heartsuit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 48 \height . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 \hfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 \hline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \hom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \hookleftarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \hookrightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \hrulefill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \hspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 \huge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \Huge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \hyphenation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 \i (dotless i) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \iff . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \Im. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \imath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 \includeonly . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70 \indent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 \index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 \indexentry . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72

\inf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\infty . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\input . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 70
\int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\intextsep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18
\iota . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\it . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\item . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 16, 19
\itemindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
\itemsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
\itshape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\j (dotless j) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\jmath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\Join . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\kappa . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\ker . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54
\kill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27
\l (␣l) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\L ( ␣L). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13
\labelenumi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\labelenumii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\labelenumiii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\labelenumiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16
\labelitemi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\labelitemii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\labelitemiii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\labelitemiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\labelsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
\labelwidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20
\lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\Lambda . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\land . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\langle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\large . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\Large . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\LARGE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\LaTeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\lbrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\lbrack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\lceil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\ldots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\le. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\leadsto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

                                                \left delim1 ... \right delim2 . . . . . . . . . . . . . . 56

\leftarrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\Leftarrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\lefteqn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
\leftharpoondown. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\leftharpoonup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49
\leftmargin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmargini . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmarginii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmarginiii . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmarginiv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmarginv . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftmarginvi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19
\leftrightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

\Leftrightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \leq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \lfloor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 \lg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \lhd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \lim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \liminf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \limsup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \line . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 \linebreak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 \linethickness . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 \linewidth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \listoffigures . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \listoftables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \listparindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \ll. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \ln. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \lnot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \location . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 \log . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \longleftarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \longleftrightarrow . . . . . . . . . . . . . . . . . . . . . . . . . 50 \longmapsto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \longrightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \lor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \lq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \makebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 \makebox (picture) . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 \makeglossary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \makeindex. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 72 \makelabels . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 \maketitle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \mapsto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \marginpar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 \marginparpush . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 \marginparsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 \marginparwidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 \markboth{left}{right} . . . . . . . . . . . . . . . . . . . . . 60 \markright{right} . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 \mathbf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \mathcal . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \mathnormal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \mathrm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \mathsf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \mathtt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \mathversion . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \max . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \mbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64 \mdseries . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \medskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \medskipamount . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \mho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \mid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \min . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \models . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \month . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 \mp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 \mu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50

\multicolumn . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29
\multiput . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24
\nabla . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\natural . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\ne. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\nearrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\neg . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\neq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\newcommand . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35
\newcounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35
\newenvironment . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36
\newfont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37
\newlength. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 35
\newline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41
\NEWLINE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
\newpage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43
\newsavebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36
\newtheorem . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 36
\ni. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\nocite . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 31
\nofiles . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71
\noindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44
\nolinebreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42
\nonumber . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17
\nopagebreak . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43
\normalfont. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\normalmarginpar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44
\normalsize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\not . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\notin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\nu. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\nwarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\o (ø). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\O (Ø). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\obeycr . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41
\oddsidemargin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4
\odot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\oe (œ) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\OE (Œ) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68
\oint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\omega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\Omega . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\ominus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50
\onecolumn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10
\opening . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\oplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
\oslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
\otimes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
\oval . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24
\overbrace{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
\overline{text}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
\owns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51
\P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\pagebreak. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 43
\pagenumbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59
\pageref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13
\pagestyle. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59

\paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 \parallel . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \parbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 \parindent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 44 \parsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \parskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 \parskip example . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \part . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 \partial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \partopsep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \perp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \phi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \pi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \Pi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \pm. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \pmod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \poptabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 \pounds . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \Pr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \prec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \preceq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \prime . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \prod . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \propto . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \protect . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37 \ps. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74 \psi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \Psi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \pushtabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 \put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 \raggedbottom . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \raggedleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 \raggedright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \raisebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65 \rangle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rbrace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rbrack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rceil . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \Re. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \ref . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 \refstepcounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 \renewenvironment. . . . . . . . . . . . . . . . . . . . . . . . . . . . 36 \restorecr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 \reversemarginpar. . . . . . . . . . . . . . . . . . . . . . . . . . . . 44 \rfloor . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rhd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \right . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \rightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \Rightarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rightharpoondown. . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rightharpoonup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 \rightleftharpoons . . . . . . . . . . . . . . . . . . . . . . . . . . 52 \rightmargin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \rm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \rmfamily . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \roman . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \rq. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67

\rule . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69
\S . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67
\savebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 65
\sbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66
\sc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\scriptsize. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\scshape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\searrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\section . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12
\selectfont. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9
\setcounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
\setlength. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
\setminus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\settodepth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
\settoheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
\settowidth . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40
\sf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\sffamily . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\sharp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\shortstack . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25
\sigma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\Sigma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\signature. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\sim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\simeq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\sinh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55
\sl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\slshape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6
\small . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7
\smallint . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\smallskip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
\smallskipamount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62
\smile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\SPACE . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61
\spadesuit. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqcap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqcup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqrt[root]{arg} . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56
\sqsubset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqsubseteq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqsupset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\sqsupseteq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\ss (ß) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68

                                                \stackrel{text}{relation} . . . . . . . . . . . . . . . . . . 56

\star . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\startbreaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\stepcounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39
\stop . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 77
\stopbreaks . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74
\subparagraph . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12
\subsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12
\subset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\subseteq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\subsubsection . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12
\succ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52
\succeq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52

\sum . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 \sup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \supset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 \supseteq . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 \surd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 52 \swarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \symbol . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \t (tie-after accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \TAB . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 \tabbingsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 \tabcolsep. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 \tableofcontents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 71 \tan . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \tanh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \tau . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \telephone. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 75 \TeX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 67 \textbf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textfloatsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \textfraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \textheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \textit . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textmd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textnormal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textrm . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textsc . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textsf . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textsl . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \texttt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \textwidth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \thanks{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \theta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \thicklines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 \thinlines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 24 \thinspace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 61 \thispagestyle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 60 \tilde . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \times . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \tiny . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \title{text} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 59 \to. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \today . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 69 \top . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \topfraction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 \topmargin. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \topsep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 \topskip . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 \totalheight . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 \triangle . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \triangleleft . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \triangleright . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \tt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 \ttfamily . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \twocolumn. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 \typein . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 \typeout . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 76 \u (breve accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \unboldmath . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 \underbrace{math}. . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \underline{text}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 56 \unitlength . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 \unlhd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \unrhd . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \uparrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \Uparrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \updownarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \Updownarrow . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \uplus . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \upshape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 \upsilon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \Upsilon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \usebox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 66 \usecounter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \usefont . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 \usepackage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 \v (breve accent) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 68 \value . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 38 \varepsilon . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \varphi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \varpi . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \varrho . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \varsigma . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \vartheta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \vdash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \vdots . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 57 \vec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \vector . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 \vee . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \verb . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32 \vert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \Vert . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 53 \vfill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 62 \vline . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 \vspace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 63 \wedge . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \widehat . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 55 \width . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 40 \wp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \wr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \xi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \Xi. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54 \year . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 39 \zeta . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 54

1

10pt option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 11pt option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 12pt option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

A

a4paper option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 a5paper option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 abstract environment . . . . . . . . . . . . . . . . . . . . . . . . . 14 array environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 article class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

B

b5paper option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 book class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

C

center environment . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

D

description environment . . . . . . . . . . . . . . . . . . . . . 15 displaymath environment . . . . . . . . . . . . . . . . . . 15, 46 document environment . . . . . . . . . . . . . . . . . . . . . . . . . 16 draft option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

E

enumerate environment . . . . . . . . . . . . . . . . . . . . . . . . 16 eqnarray environment . . . . . . . . . . . . . . . . . . . . . . . . . 17 equation environment . . . . . . . . . . . . . . . . . . . . . 17, 46 executivepaper option . . . . . . . . . . . . . . . . . . . . . . . . . 4

F

figure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 final option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 fleqn option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 flushleft environment . . . . . . . . . . . . . . . . . . . . . . . . 18 flushright environment. . . . . . . . . . . . . . . . . . . . . . . 19

I

itemize environment . . . . . . . . . . . . . . . . . . . . . . . . . . 19

L

landscape option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 legalpaper option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 leqno option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 letter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 letter class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 letterpaper option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 list . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 lR box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 lrbox . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 64

M

math environment . . . . . . . . . . . . . . . . . . . . . . . . . . 21, 46 minipage environment . . . . . . . . . . . . . . . . . . . . . . . . . 21

N

notitlepage option. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

O

onecolumn option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 oneside option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 openany option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 openbib option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 openright option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

picture . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22

Q

quotation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 quote . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

R

report class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

S

slides class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

T

tabbing environment . . . . . . . . . . . . . . . . . . . . . . . . . . 26 table . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 tabular environment . . . . . . . . . . . . . . . . . . . . . . . . . . 28 thebibliography . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30 theorem environment . . . . . . . . . . . . . . . . . . . . . . . . . . 31 titlepage environment . . . . . . . . . . . . . . . . . . . . . . . . 32 titlepage option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 twocolumn option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 twoside option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

V

verbatim environment . . . . . . . . . . . . . . . . . . . . . . . . . 32 verse environment . . . . . . . . . . . . . . . . . . . . . . . . . . . . 32